### v1

# SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

Abstract
Extreme low-bit quantization is critical for efficiently deploying Large Language Models (LLMs), yet it often leads to severe performance degradation at 2-bits and even 4-bits (e.g., MXFP4). We present SignRoundV2, a post-training quantization framework that is highly effective even without mixed-precision. SignRoundV2 introduces (1) a fast sensitivity metric that combines gradient information with quantization-induced deviations to guide layer-wise bit allocation, and (2) a lightweight pre-tuning search for quantization scales to improve extremely low-bit quantization. These components allow SignRoundV2 to close the gap with full-precision models. Extensive experiments indicate that our method sustains competitive accuracy for LLMs, achieving production-grade performance with $\sim$1% variance at $4\text{-}5$ bits and strong results even at 2 bits. The implementation is available at https://github.com/intel/auto-round.
SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs
Wenhua Chengthanks: Correspondence: wenhua.cheng@intel.com  and Weiwei Zhangthanks: These authors contributed equally  and Heng Guo22footnotemark: 2  and Haihao Shen Intel
1 Introduction
The advent of Large Language Models (LLMs) such as GPT (brown2020language), LLaMA (touvron2023llama; meta2024llama3), Qwen (bai2023qwen; yang2025qwen3) and DeepSeek (liu2024deepseek) has marked a major shift in modern artificial intelligence. As parameter counts scale from billions to hundreds of billions, these models have demonstrated unprecedented capabilities in reasoning, coding, multimodal understanding, and autonomous agent behaviors. However, such scaling comes at the cost of dramatically increased memory consumption, bandwidth pressure, and inference latency, posing significant barriers for real-world deployment, especially on resource-constrained environments such as consumer GPUs, CPUs, or edge devices.
To alleviate these limitations, post-training quantization (PTQ) has emerged as one of the most practical approaches because it avoids costly re-training and can be applied to a wide range of pretrained LLMs (xiao2023smoothquant; frantar2022gptq; lin2024awq; cheng2024optimize). By compressing weights and activations to low-bit representations, PTQ can yield substantial reductions in memory footprint and hardware cost. In particular, the push toward extreme low-bit quantization, for example, sub-4-bit weight-only quantization and 4-bit activation quantization, has become a crucial enabler for democratizing LLM deployment. However, achieving this while preserving accuracy remains highly challenging due to increased quantization error, layer sensitivity imbalance, and the lack of hardware-friendly data types.
Figure 1: Average accuracy of pure 2-bit (W2A16) models on Llama 2/3 70B. See detailed results in Table 1.
Recent research has moved beyond conventional W4A16 and W8A8 schemes (where WA denotes weight and activation bit-widths) toward more expressive low-bit data types. For example, MXFP4 (mxfp4) and NVFP4 (nvpf4) are newly introduced floating-point variants optimized for modern accelerators, demonstrating that well-designed data types can maintain reasonable accuracy even under aggressive activation quantization. In parallel, weight-only quantization continues to push the limits of bit efficiency. In particular, BitNet (wang2025bitnet; wu2025bitnet1) explores quantization down to approximately 1.58 bits, showing that extremely low-bit representations can still be learned if specialized training or distillation procedures are used. However, these methods require training from scratch or costly post-training distillation, making them less suitable for scenarios where fast, scalable, and hardware-efficient PTQ is desired.
A natural direction for overcoming the limitations mentioned above is Adaptive Bit-Width Quantization (ABQ). ABQ mitigates the shortcomings of uniform-bit quantization by allocating higher precision to more sensitive layers while aggressively compressing more robust ones (zhou2018adaptive; dong2019hawq; wang2019haq; NEURIPS2024_2c30a37c). The central challenge therefore becomes the design of an accurate and efficient layer sensitivity metric to guide bit allocation. Early methods estimate sensitivity using the Hessian-related matrix (dong2019hawq; NEURIPS2024_2c30a37c), or rely on layer-wise/model-wise output distortion measured on a small calibration dataset (hubara2020improving). However, these heuristics often exhibit imperfect correlation when applied layer-wise, and model-wise approaches typically require slow calibration, especially for large language models (LLMs). In contrast, learning-based or reinforcement learning based approaches automate precision assignment (lou2019autoq; wang2019haq), but at the cost of substantial compute overhead, contradicting the primary objective of efficient PTQ.
To address these challenges, we introduce SignRoundV2, an enhanced post-training quantization framework building upon the original SignRound (cheng2024optimize). Our method is driven by two key insights. First, an effective sensitivity metric must capture both the local parameter distortion caused by quantization and its global impact on task loss. We design a simple yet highly effective metric that integrates gradient information with quantization-induced parameter deviation, yielding a far more reliable signal for ABQ than existing proxies. Then given a target average bit budget, SignRoundV2 employs dynamic programming to convert layer-wise sensitivity scores into an optimal bit configuration. Second, we observe that the initialization of quantization parameters especially scale is critical for extremely low-bit stability, consistent with findings from recent PTQ literature (lin2024awq; ma2024era). Accordingly, inspired by the importance matrix in llama.cpp (ggerganov_llamacpp), we introduce a lightweight pre-tuning search that rapidly identifies high-quality initialization values before the main tuning stage, substantially improving recoverable accuracy. Figure 1 compares our method with state-of-the-art baselines, showing that it achieves similar performance at significantly lower cost and avoids the issues associated with Quantization Aware Training (QAT), which are discussed in Section 2. In summary, our contributions are three-fold.
  • We propose a simple and efficient sensitivity metric that integrates gradient information with quantization-induced parameter deviations. This metric captures both local parameter distortions and their global impact on task loss, providing a reliable guide for adaptive bit-width quantization.
  • We incorporate a lightweight pre-tuning search for quantization scales prior to the main tuning stage, enhancing both stability and final model accuracy under extremely low-bit settings. Experimental results demonstrate that this strategy achieves performance on par with recent efficient QAT methods.
  • We develop a fast, adaptive mixed-precision quantization method and show that it achieves competitive performance compared to existing QAT approaches, while incurring substantially lower quantization cost.
2 Related Work
Quantization-Aware Training (QAT).
QAT integrates learnable quantizers directly into the training loop, minimizing the task loss and fine-tuning additional parameters beyond the quantizers introduced by low-precision operations, as described in EfficientQAT (chen2025efficientqat), LLM-QAT (liu2023llm), DL-QAT (ke2024dl), and BitDistiller (du2024bitdistiller). Recent work(chee2023quip; tseng2024quip; egiazarian2024extreme) has revisited vector quantization for extremely low-bit settings, but these methods still rely on QAT to achieve high accuracy. Despite their effectiveness, QAT methods suffer from several practical limitations. First, they optimize the task loss or the unquantized parameters of the model, which can lead to catastrophic forgetting (luo2025empirical; kalajdzievski2024scaling), particularly in domains not covered by fine-tuning. Second, QAT requires careful hyperparameter tuning, and its actual quantization cost is often much higher than what is reported, typically far exceeding that of post-training quantization. Third, QAT generally requires substantially more data than PTQ to mitigate overfitting, further increasing the overall resource demands.
Post-Training Quantization (PTQ).
Post-training quantization (PTQ) avoids additional fine-tuning of non-quantized parameters and task loss, and therefore offers a simple and resource-efficient pipeline for compressing LLMs (nagel2019data; liu2021post; frantar2022optimal; yao2021hawq; xiao2023smoothquant; wei2023outlier; dettmers2022gpt3). Because LLM decoding is typically memory-bound, weight-only PTQ has become the dominant approach (frantar2022gptq; lin2024awq; cheng2023teq; yao2024exploring; badri2023hqq; kim2023squeezellm; cheng2024optimize) in LLM inference. However, existing weight-only methods still suffer from a significant accuracy drop at extremely low bit-widths (e.g., 2-bit).
Mixed Precision Quantization.
Mixed-precision quantization assigns heterogeneous bit-widths across layers or channels to better align with their sensitivity. Early approaches include Hessian-based methods such as HAWQ (dong2019hawq) and BAQ (zhang2025baq), as well as reinforcement-learning–based schemes such as HAQ (wang2019haq). However, these methods are prohibitively expensive for multi-billion-parameter LLMs due to second-order matrix computations or extensive policy evaluations. More recent systems, such as MicroMix (liu2025micromix), apply FP4/FP6/FP8 at the channel level via MXFP microscaling kernels, but their reliance on specialized hardware limits the portability. In practice, Llama.cpp (ggerganov_llamacpp) adopts a more lightweight strategy that relies on heuristic rules to assign mixed precisions for each model and quantization scheme. Recent works (duanmu2025mxmoe; tao2025moqae; chitty2025mopeq) focus on mixed-precision settings for mixture-of-expert (MoE) models, which have become particularly popular for large-scale models.
Rounding and Optimized Quantizer Search.
Rounding critically affects the quality of quantized weights. AdaRound (nagel2020up) formulates weight rounding as a per-weight binary optimization problem, where a second-order Taylor expansion of the task loss approximates quantization-induced perturbations, and a layer-wise local reconstruction loss is minimized via continuous relaxation. This formulation has inspired subsequent works such as BRECQ (li2021brecq), SignRound (cheng2024optimize). FlexRound (lee2023flexround) increases rounding flexibility through element-wise scaling, albeit at the cost of introducing numerous hyperparameters. Oscillation-Free training (liu2023oscillation) highlights instability arising from learnable rounding parameters. Activation-dependent rounding in AQuant (li2022efficient) reduces activation error but is not applicable to weight-only inference. SignRound (cheng2024optimize) jointly optimizes rounding and weight clipping using sign-based gradient descent, achieving substantial improvements, particularly at extremely low bit-widths such as 2 bits.
3 Methodology
This section presents our methodology. We begin in Section 3.1 by introducing the sensitivity metric, DeltaLoss, which guides the mixed-bit assignment. Section 3.2 then leverages DeltaLoss together with dynamic programming to generate adaptive bit-widths for each layer. In Section 3.3, we describe our pre-tuning search for quantization parameters. Finally, Section 3.4 lists all major hyperparameters and details a practical trick for loss calculation.
Before introducing our method, we first provide a brief overview of quantization and SignRoundV1  (cheng2024optimize). The following operation can be used to quantize and dequantize (qdq) the weights $\mathbf{W}$:
$\mathbf{qdq(W)}=s\cdot\text{clip}\Big(\Big\lfloor\frac{\mathbf{W}}{s}\Big\rceil,n,m\Big),\quad n,m\in\mathbb{N},$ (1)
where the rounding operation $\lfloor\cdot\rceil$ is typically performed using the Round-to-Nearest (RTN) method. The scale factor $s$ is defined as:
$s=\frac{\max(\mathbf{W})-\min(\mathbf{W})}{2^{\text{bit}}-1},$ (2)
and for simplicity, we ignore the zero point.
SignRoundV1 (cheng2024optimize) introduces three trainable parameters $v$, $\alpha$ and $\beta$, to reduce the quantization error. The parameter $v$ enhances the rounding operation
$\mathbf{qdq(W)}=s\cdot\text{clip}\Big(\Big\lfloor\frac{\mathbf{W}}{s}+v\Big\rceil,n,m\Big),\quad n,m\in\mathbb{N},$ (3)
while $\alpha$ and $\beta$ refine the scale and zero-point through
$s=\frac{\max(\mathbf{W})*\alpha-\min(\mathbf{W})*\beta}{2^{\text{bit}}-1}$ (4)
(a) W2A16 layer-wise deltaloss sensitivity
(b) MXFP4 layer-wise deltaloss sensitivity
Figure 2: Layer-wise DeltaLoss sensitivity of Llama-3.1-8B-Instruct under W2A16 and MXFP4.
3.1 DeltaLoss Sensitivity Metric
Prior work  (dong2019hawq; NEURIPS2024_2c30a37c) largely ignores first-order gradient information and instead relies on second-order approximations, such as the Hessian or Fisher information matrix, to estimate sensitivity, under the assumption that gradients are typically near zero. However, this assumption breaks down when quantization induces a large loss. Moreover, although second-order metrics capture local curvature, they share a fundamental limitation with purely gradient-based methods: both fail to reflect the actual impact of quantization on the loss, providing only a relative trend rather than a reliable measure of true loss.
To address this limitation, we employ a first-order Taylor expansion to directly estimate the loss change induced by quantization, using the commonly adopted cross-entropy loss. Specifically, for a given layer with weight tensor $\mathbf{W}$ and activation tensor $\mathbf{A}$, the loss variation caused by quantizing the layer can be approximated using the first-order Taylor expansion:
$\Delta\mathcal{L}\approx g_{aq}\circ(A_{\text{f}}-A_{\text{q}})+g_{wq}\circ(W_{\text{f}}-W_{\text{q}})$ (5)
where:
  • $g_{aq}=\frac{\partial\mathcal{L}}{\partial A_{q}}$ represents the gradient of the loss with respect to the qdq activation tensor
  • $g_{wq}=\frac{\partial\mathcal{L}}{\partial W_{q}}$ represents the gradient of the loss with respect to the qdq weight tensor
  • $A_{\text{f}}$ and $W_{\text{f}}$ denote the full-precision activation and weight values, respectively
  • $A_{\text{q}}$ and $W_{\text{q}}$ denote the qdq activation and weight values, respectively
  • $\circ$ denotes the Hadamard product
In practice, we found that using the absolute value function yields more robust results, so we redefine the metric as
$\Delta\mathcal{L}\approx\left|(g_{aq}\circ(A_{\text{f}}-A_{\text{q}})\right|+\left|g_{wq}\circ(W_{\text{f}}-W_{\text{q}})\right|$ (6)
This formulation effectively combines gradient information and parameter distortion to yield a more accurate and holistic sensitivity estimate for each layer.
To reduce memory consumption, activation quantization has been identified as the dominant contributor to quantization loss as shown in (xiao2023smoothquant; wei2023outlier). Therefore, within the context of joint weight–activation quantization, we primarily focus on activation-induced distortion and omit the weight quantization term in our sensitivity formulation.
$\Delta\mathcal{L}$ $\approx\left|g_{aq}\circ(A_{\text{f}}-A_{\text{q}})\right|$ (7)
Figure 2 visualizes the layer-wise sensitivity measured by DeltaLoss. It is evident that sensitivity varies across different data types, making it prohibitively labor-intensive to rely on heuristic rules to manually set quantization parameters for each scheme, as done in llama.cpp (ggerganov_llamacpp). Moreover, certain layers within a block, such as down_proj, exhibit higher sensitivity, suggesting that traditional block-wise fallback strategies may be suboptimal.
3.2 Layer-wise Bit Allocation
We formulate layer-wise bit allocation as a discrete optimization problem. Let $n$ denote the total number of layers, and let $B=\{b_{1},\dots,b_{K}\}$ be the set of allowed bit-widths. For each layer $i\in\{1,\dots,n\}$ and bit-width $b\in B$, we introduce a binary variable $I_{i,b}\in\{0,1\}$, where $I_{i,b}=1$ indicates that layer $i$ is assigned bits $b$.
For simplicity, we assume identical bit-widths for both weights and activations, although this formulation can be extended to accommodate different bit-widths or data types across layers.
Let $\Delta L_{i}(b)$ denote the delta loss when quantizing layer $i$ to $b$ bits. Given a target average bit-width $T$, the optimization problem can be written as:
$\min_{I_{i,b}}$ $\sum_{i=1}^{n}\sum_{b\in B}\Delta L_{i}(b)\,I_{i,b}$ (8)
s.t. $\sum_{b\in B}I_{i,b}=1,\quad\forall i=1,\dots,n,$
$\sum_{i=1}^{n}\sum_{b\in B}b\,I_{i,b}*P_{i}\leq T*\sum_{i}{P_{i}},$
$I_{i,b}\in\{0,1\},\quad\forall i=1,\dots,n,\;b\in B.$
This optimization problem can be solved using dynamic programming  (bellman1966dynamic) or as an Integer Linear Program  (wolsey2020integer; bertsimas1997introduction), both of which are well-established methods. We omit the details here and refer the reader to the existing literature.
3.3 Quantization Parameters Initialization
In SignRoundV1 (cheng2024optimize), all quantization parameters are initialized with trivial values, i.e., weight clipping is set to 1.0 and the rounding perturbation is 0. A better initialization could potentially improve performance, similar to the effect of careful initialization in training neural networks  (he2015delving; glorot2010understanding). Although directly initializing the rounding values is challenging, prior works (lin2024awq; badri2023hqq; ggerganov_llamacpp) have proposed training-free methods to search for weight clipping by aligning layer outputs or optimizing certain objectives. While effective, these approaches require layer-wise forward passes, which can be slow, or rely on floating-point zero points for better accuracy.
Inspired primarily by the importance matrix in llama.cpp (ggerganov_llamacpp), we adopt a simple variant to initialize the scale. The search process is formulated as
$\min_{s\in S}\;\frac{1}{N}\sum_{i=1}^{N}\left((W_{F}-W_{q})\circ max(\bar{A})^{2}\right)_{i}^{2}$ (9)
where $S$ denotes a predefined set of candidate scales, and $max(\bar{A})$ represents the input channel-wise maximum absolute values, calibrated from a set of samples.
For symmetric quantization, which is the default setting used in this work, the set $S(s_{0},...,s_{c})$ is constructed from the following candidates
$s_{i}=\frac{\max(\mathbf{|W|})}{2^{\text{bit}-1}+\epsilon_{i}}.$ (10)
with $\epsilon_{i}$ sampled from
$\epsilon\in\mathrm{range}(-0.9,\,0.9,\,\text{step}=0.01)$ (11)
Following SignRoundV1 (cheng2024optimize), the best scale $s_{init}$ that minimizes Eq. 9 is further refined by a learnable parameter $\alpha$, which we constrain to the range [0.5, 1.5] in our experiments.
$s=s_{init}*\alpha$ (12)
3.4 Hyperparameters and Tricks
For DeltaLoss, we use only 16 calibration samples with a sequence length of 256. For tuning, we follow the setup in SignRoundV1 (cheng2024optimize). Each transformer block is optimized for 200 steps using sign gradient descent, with a learning rate of $1/steps$ and a batch size of 8. The sequence length is fixed at 2048. To reduce quantization cost, we lower the default number of calibration samples from 512 to 128. We additionally provide a higher-cost recipe for improved accuracy, denoted as Ours*, where the number of steps is increased from 200 to 500, the learning rate is set to $2.0/steps$ and the number of calibration samples is restored to 512. Automatic mixed precision (AMP) is applied throughout to improve tuning efficiency.
To enhance tuning stability, we exclude the top-$k$ largest losses in a batch when computing the mean squared error between the quantized block output and the full-precision block, where $k$ is set to 0.1% of the total number of elements. Let $\mathrm{TopK}(\mathcal{L},k)$ denote the set of the $k$ largest losses in $\mathcal{L}$. The effective loss used for optimization is defined as follows, where $\setminus$” denotes the exclusion operation:
$L_{\text{eff}}=\sum_{L_{i}\in\mathcal{L}\setminus\text{TopK}(\mathcal{L},k)}L_{i}.$ (13)
Table 1: Average accuracies across five tasks (Section 4.2) using W2A16 at 2 bits and mixed W2A16 and W4A16G128 at 2.5 bits. "G" denotes the group size, and "Avg. Bits" denotes the average per-weight bit-width. Ours* indicates our improved recipe (Section 3.4). ’-’ denotes the missing data from the original work.
Avg. bits Method Group Size Llama2-7B Llama2-13B Llama2-70B Llama3-8B Llama3-70B
16 16bits - 64.66 67.44 72.71 68.64 75.28
2 AQLM 2x8 57.61 62.22 69.85 - -
AQLM 1x16 61.85 64.95 70.84 64.10 70.10
QuIP# - 60.61 64.44 70.91 - -
EfficientQAT 128 59.50 63.88 68.93 59.37 67.57
EfficientQAT 64 60.14 64.48 69.48 60.76 67.89
GPTQ 128 41.56 48.29 34.38 - -
AWQ 128 34.74 35.99 35.49 - -
OmniQ 128 46.98 53.56 54.87 52.66 60.06
SRV1 128 54.50 60.72 67.70 55.25 64.76
Ours 128 57.88 61.88 68.39 59.02 69.02
Ours* 128 58.67 62.34 68.82 59.97 70.16
Ours* 64 59.04 62.81 69.30 60.38 70.93
2.5 Ours 128 59.96 63.65 70.60 62.76 72.37
Ours* 128 60.44 63.73 70.04 62.60 72.68
Ours 64 60.08 64.09 70.25 63.67 71.97
Ours* 64 60.28 64.57 70.60 63.56 72.32
Table 2: The average accuracies and recovery rates (%) across 10 tasks(detailed in Section 4.3) with MXFP4 at 4bits and mixed MXFP4/8 for others. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning. The "I" in the model name is short for "Instruct".
Avg. Bits Method Llama3.1-8B-I Llama3.1-70B-I Qwen2.5-7B-I Qwen3-8B Qwen3-32B
16 16bits 64.16 70.00 65.67 63.24 67.00
4 RTN 58.31 (90.88%) 68.71 (98.16%) 60.62 (92.32%) 58.54 (92.57%) 65.07 (97.12%)
SRV1 60.72 (94.64%) 69.01 (98.60%) 64.06 (97.55%) 60.25 (95.28%) 66.92 (99.88%)
Ours 61.34 (95.59%) 69.32 (99.04%) 63.37 (96.50%) 61.89 (97.87%) 66.86 (99.79%)
4.5 DL 60.64 (94.50%) 69.78 (99.70%) 63.13 (96.13%) 59.81 (94.58%) 66.05 (98.58%)
Ours 62.33 (97.14%) 69.96 (99.95%) 64.51 (98.24%) 62.28 (98.49%) 66.89 (99.83%)
5 DL 62.57 (97.51%) 70.10 (100.15%) 64.04 (97.52%) 61.45 (97.18%) 66.00 (98.51%)
Ours 63.19 (98.49%) 70.31 (100.45%) 65.04 (99.03%) 62.54 (98.89%) 67.17 (100.25%)
6 DL 63.64 (99.18%) 70.56 (100.80%) 65.45 (99.67%) 62.00 (98.04%) 66.61 (99.42%)
Ours 64.12 (99.93%) 70.50 (100.71%) 65.30 (99.43%) 62.19 (98.34%) 67.21 (100.32%)
Table 3: Comparison with head-layer (Head)/ tail-layer (Tail) heuristics at 4.5-bit and 5-bit MXFP with only DL, without tuning. The average accuracies are across 10 tasks (detailed in Section 4.3). The "I" in the model name is short for "Instruct".
Avg. bits Method Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
4.5 Head 8-bit 59.18 60.63 58.92
Tail 8-bit 58.93 61.38 59.26
DL 60.64 63.13 59.81
5 Head 8-bit 60.82 61.75 59.43
Tail 8-bit 60.12 61.35 59.65
DL 62.57 64.04 61.45
6 Head 8-bit 59.88 62.22 60.54
Tail 8-bit 62.65 63.39 61.53
DL 63.64 65.45 62.00
Table 4: The average accuracies across 10 tasks(Section 4.3) at an average of 3 bits (W2G128 / W4G128). The "I" in the model name is short for "Instruct".
Avg. bits = 3 Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
Head 4-bit 31.98 32.70 31.96
Tail 4-bit 60.58 37.98 45.36
Ours 61.48 40.58 48.62
4 Experiments
4.1 Experimental Setup
Models and Benchmarks.
We evaluate SignRoundV2 on two major LLM families, LLaMA (touvron2023llamav2; grattafiori2024llama) and Qwen (bai2023qwen; yang2025qwen3), under both low-bit and mixed-precision quantization settings across standard benchmarks. All evaluations follow the lm-eval-harness framework (eval-harness). The benchmark suite includes ARC-Challenge (clark2018think), ARC-Easy (clark2018think), BoolQ (clark2019boolq), HellaSwag (Zellers_2019), LAMBADA (paperno2016lambada), MMLU (hendrycks2020measuring), OpenBookQA (mihaylov2018can), PIQA (bisk2020piqa), TruthfulQA (lin2022truthfulqa), and WinoGrande (sakaguchi2021winogrande).
Quantization Configurations.
Following SignRoundV1 (cheng2024optimize), we use The Pile (gao2020pile) as the calibration dataset and adopt symmetric quantization. All experiments are conducted on NVIDIA A100 80GB GPUs. For simplicity, we omit the bit costs associated with scales and zero points. For MXFP4 and MXFP8, we follow the standard definitions in (mxfp4).
4.2 Comparison of Results at W2A16/W4A16
Table 1 presents the average accuracies across 5 tasks (ARC-Challenge, ARC-Easy, HellaSwag, PIQA and WinoGrande) for W2A16 with and without mixed-precision quantization on various LLaMA models. At pure W2A16 setting, our method consistently outperforms existing PTQ approaches such as GPTQ  (frantar2022gptq), AWQ  (lin2024awq), OmniQuant  (shao2023omniquant), and SignRoundV1  (cheng2024optimize) by a substantial margin. Compared to high cost methods like AQLM  (egiazarian2024extreme), QuIP# (tseng2024quip), and EfficientQAT  (chen2025efficientqat), our approach achieves comparable performance on large models, while slightly trailing on smaller models. At slightly higher bit widths (e.g., 2.5 bits), with the support of DeltaLoss, our method continues to deliver strong accuracy, demonstrating the effectiveness of adaptive layer-wise quantization. Overall, these results confirm that our approach reliably preserves model quality under extreme low-bit mixed-precision settings. Detailed accuracy results for each task are provided in Appendix A.
4.3 Comparison of Results at MXFP4/8
Table 2 reports the average accuracies and recovery rates across 10 tasks (all tasks listed in Section 4.1) for MXFP4/8 mixed-precision quantization. Since few prior works report results for standard MXFP4, we primarily compare against our own baselines. The results show that our method consistently outperforms RTN and SignRoundV1 (SRV1) at low average bit widths (4-5 bits), achieving higher accuracy and recovery rates across nearly all evaluated models. At 4 bits, our approach improves recovery by $\sim$1-3% compared to SRV1. At slightly higher bit widths (5-6 bits), both DeltaLoss-only (DL) and our method approach full-precision accuracy, but our method achieves the highest recovery in most scenarios, reaching near or above 99% across all models. Overall, these results demonstrate that our approach effectively preserves model performance under extreme low-bit mixed-precision quantization while minimizing accuracy loss. Detailed accuracy results for each task are provided in Appendix A.
4.4 Comparison of Mixed Bits Recipes
The results in Tables 3 and 4 demonstrate that our method consistently outperforms simple head-layer and tail-layer heuristics across a range of precision budgets and model architectures. For MXFP quantization (4.5, 5, and 6 bits), the DL strategy achieves the highest average accuracy for all three models without requiring any tuning, showing clear improvements over allocating 8-bit precision only to the head (near the LM head) or tail layers (near the embedding). At lower precision (average 3 bits using W2G128/W4G128), the performance gap becomes even more obvious: both head- and tail-focused heuristics degrade substantially, while our method remains robust and delivers the best accuracy across all evaluated models.
4.5 Ablation Study of Initialization
We evaluate the impact of our initialization strategy on Qwen3-8B and Llama3.1-8B-Instruct under the W2A16G64 setting on SignRoundV1 with or without initialization. As shown in Table 5, enabling initialization consistently improves accuracy across all five tasks. Overall, the results demonstrate that proper pre-tuning initialization significantly stabilizes optimization and leads to substantial accuracy improvements for both models.
Table 5: Ablation study of pre-tuning initialization for Qwen3-8B and Llama3.1-8B-Instruct at W2A16G64.
Model Method Arc-c hella. lamb. mmlu wino.
Qwen3-8B W/O init 34.90 44.87 47.18 54.09 63.85
With init 43.69 45.12 54.61 56.12 66.22
Llama3.1-8B-I W/O init 36.43 46.20 52.57 48.05 65.67
With init 36.43 47.32 60.41 50.11 67.40
4.6 Quantization Cost
Table 6 compares the end-to-end quantization time across several representative methods. Traditional QAT-based and vector quantization approaches such as EfficientQAT (chen2025efficientqat), QuIP#  (tseng2024quip), and AQLM  (egiazarian2024extreme) require tens to hundreds of hours to converge. Besides, we argue that QAT requires careful hyperparameter tuning, and its actual quantization cost is often much higher than what is reported. In contrast, SignRoundV1 completes in only 2.2 hours. Our method maintains a similarly low computational cost, requiring 2.5 hours, while the enhanced version (Ours*) takes 6 hours due to additional tuning samples and steps. Overall, our approach achieves competitive accuracy with substantially lower time overhead compared with prior methods.
Table 6: Time Cost Comparison for Llama2-70B
Method One A100-80GB? GPU hours
SignRoundV1 2.2
EfficientQAT 41
QuIP# × 270
AQLM 336
Ours 2.5
Ours* 6
Table 7 summarizes the VRAM and runtime overhead of computing DeltaLoss for weight-only quantization (e.g., W2A16), with MXFP4 exhibiting comparable cost. The results show that DeltaLoss remains computationally practical even for 70B-scale models.
Table 7: VRAM and Time Cost of DeltaLoss For W2A16
Model VRAM Cost (GB) Time Cost (s)
Qwen3-8B 15 $60\times len\ of\ options$
Qwen3-32B 30 $180\times len\ of\ options$
Llama-2-70B 40 $420\times len\ of\ options$
5 Conclusion
We introduced SignRoundV2, a practical framework for extreme low-bit post-training quantization of LLMs. By combining a gradient-informed sensitivity metric with a lightweight pre-tuning search, SignRoundV2 enables stable and accurate quantization even at aggressive bit widths. Experiments show that it delivers production-grade accuracy at 4–5 bits and strong performance at 2 bits. Our results demonstrate that effective PTQ can significantly reduce the deployment cost of LLMs while preserving model quality.
6 Limitations
Despite its strong performance, SignRoundV2 has several limitations. First, without mixed precision, it still exhibits a noticeable gap from the full-precision model at extremely low bit widths, particularly for smaller models. Second, the bit configuration is fixed before tuning, without considering how much quantization error could be mitigated during the tuning process. Third, the method requires gradient computation, which limits its applicability in frameworks such as ONNX that do not support gradient-based optimization.
Ethics Statement
Our research aims to advance knowledge in LLM quantization. SignRoundV2 utilizes open-source models and publicly available datasets, and is not tied to particular applications, requiring only minimal fine-tuning steps on the original models. This ensures that the technical details of our method carry no potential ethical implications. We acknowledge the contributions of the creators and maintainers of these resources and provide citations to the original sources.
Appendix A Detailed Data
This appendix provides detailed accuracy results for each task.
Table 8: The detailed accuracies of 5 tasks for Llama2 models with INT2/4 mixed precision. "Avg. Bits" indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. Piqa Wino. Avg.
Llama-2-7B 16Bits 16 - 42.83 75.97 57.27 77.91 69.30 64.66
AQLM 2 2x8 32.85 66.92 49.96 73.07 65.27 57.61
AQLM 2 1x16 39.68 74.07 53.42 76.88 65.19 61.85
QuIP# 2 - 37.88 71.84 52.19 75.46 65.67 60.61
EfficientQAT 2 128 36.52 69.78 50.84 74.16 66.22 59.50
EfficientQAT 2 64 36.86 70.96 51.58 75.30 65.98 60.14
GPTQ 2 128 21.25 40.45 32.59 58.32 55.17 41.56
AWQ 2 128 21.08 24.62 25.69 52.34 49.96 34.74
OmniQ 2 128 23.46 50.13 40.28 65.13 55.88 46.98
SignRoundV1 2 128 32.25 65.99 40.28 72.96 61.01 54.50
Ours 2 128 33.53 69.32 48.17 74.43 63.93 57.88
Ours 2 64 32.93 67.97 48.54 74.21 65.59 57.85
Ours* 2 128 34.64 69.95 48.92 74.43 65.43 58.67
Ours* 2 64 35.32 69.99 49.30 74.16 66.41 59.04
Ours 2.5 128 35.41 71.46 50.62 75.30 67.01 59.96
Ours 2.5 64 35.92 71.25 51.44 75.35 66.46 60.08
Ours* 2.5 128 36.01 72.01 51.19 74.92 67.25 60.28
Ours* 2.5 64 36.09 71.72 51.85 74.81 67.72 60.44
Llama-2-13B 16Bits 16 - 47.18 78.32 60.25 79.22 72.22 67.44
AQLM 2 2x8 40.10 73.06 54.62 77.09 66.22 62.22
AQLM 2 1x16 43.52 75.25 57.62 78.29 70.09 64.95
QuIP# 2 - 42.92 75.72 56.53 77.97 69.06 64.44
EfficientQAT 2 128 42.83 75.04 55.66 76.99 68.90 63.88
EfficientQAT 2 64 41.89 74.83 55.27 77.04 68.36 63.48
GPTQ 2 128 21.93 55.60 41.06 67.08 55.80 48.29
AWQ 2 128 23.12 26.22 25.80 52.99 51.85 35.99
OmniQ 2 128 30.29 63.22 46.23 70.13 57.93 53.56
SignRoundV1 2 128 38.57 71.17 53.35 76.17 64.33 60.72
Ours 2 128 39.08 73.78 52.96 75.68 67.88 61.88
Ours 2 64 40.53 74.33 53.87 76.44 68.43 62.72
Ours* 2 128 40.44 74.24 53.15 75.90 67.96 62.34
Ours* 2 64 40.78 74.58 53.73 76.77 68.19 62.81
Ours 2.5 128 41.04 75.42 55.39 76.77 69.61 63.65
Ours 2.5 64 41.13 75.55 55.81 77.42 70.56 64.09
Ours* 2.5 128 40.02 75.08 55.61 76.82 71.11 63.73
Ours* 2.5 64 42.83 76.43 55.97 77.37 70.24 64.57
Llama-2-70B 16Bits 16 - 54.44 82.70 64.77 82.15 77.98 72.41
AQLM 2 2x8 51.45 79.76 61.94 80.47 75.61 69.85
AQLM 2 1x16 52.99 81.36 62.78 81.07 76.01 70.84
QuIP# 2 - 52.65 81.90 62.86 81.39 75.77 70.91
EfficientQAT 2 128 49.23 80.01 61.58 80.20 73.64 68.93
EfficientQAT 2 64 50.77 80.13 61.78 80.14 74.59 69.48
GPTQ 2 128 22.70 25.08 25.04 49.51 49.57 34.38
AWQ 2 128 22.35 25.76 25.46 52.50 51.38 35.49
OmniQ 2 128 33.28 67.21 35.45 74.10 64.33 54.87
SignRoundV1 2 128 46.59 78.37 59.65 79.00 74.90 67.70
Ours 2 128 49.06 79.59 59.99 78.35 74.98 68.39
Ours 2 64 49.06 78.91 60.21 79.00 75.61 68.56
Ours* 2 128 60.23 80.22 60.23 80.22 76.09 68.82
Ours* 2 64 50.60 79.92 60.52 79.22 76.24 69.30
Ours 2.5 128 51.71 80.56 61.77 79.82 76.64 70.10
Ours 2.5 64 50.77 80.77 61.91 80.30 77.51 70.25
Ours* 2.5 128 50.64 80.64 61.68 79.89 77.35 70.04
Ours* 2.5 64 51.02 81.31 62.29 80.58 77.82 70.60
Table 9: The detailed accuracies of 5 tasks for Llama3 models with INT2/4 mixed precision. "Avg. Bits" indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. Piqa Wino. Avg.
Llama-3-8B 16Bits 16 - 50.17 80.09 60.14 79.54 73.24 68.64
AQLM 2 1x16 41.21 74.24 55.44 77.80 71.82 64.10
EfficientQAT 2 128 36.01 69.15 50.74 75.30 65.67 59.37
EfficientQAT 2 64 37.03 71.17 51.86 76.03 67.72 60.76
SignRoundV1 2 128 34.22 67.09 44.15 71.65 59.12 55.25
Ours 2 128 35.67 70.03 47.70 72.96 63.14 57.90
Ours 2 64 37.37 70.29 48.69 74.59 66.14 59.42
Ours* 2 128 38.99 72.18 49.19 73.94 65.57 59.97
Ours* 2 64 39.42 71.42 50.04 74.43 66.61 60.38
Ours 2.5 128 41.72 74.87 52.81 75.57 68.82 62.76
Ours 2.5 64 43.34 75.42 53.27 76.06 70.24 63.67
Ours* 2.5 128 41.38 74.87 53.31 75.46 67.96 62.60
Ours* 2.5 64 43.00 74.87 54.13 76.28 69.53 63.56
Llama-3-70B 16Bits 16 - 60.24 86.83 66.40 82.21 80.74 75.28
AQLM 2 1x16 50.34 78.83 63.47 79.65 78.22 70.10
EfficientQAT 2 128 48.81 79.25 60.75 79.60 69.46 67.57
EfficientQAT 2 64 49.06 77.40 61.60 77.37 74.03 67.89
SignRoundV1 2 128 47.87 79.71 43.96 78.62 73.64 64.76
Ours 2 128 49.91 80.81 59.75 79.49 75.14 69.02
Ours 2 64 52.22 81.73 60.93 80.09 75.69 70.13
Ours* 2 128 52.05 82.41 60.41 79.38 76.56 70.16
Ours* 2 64 54.95 82.53 61.05 79.82 76.32 70.93
Ours 2.5 128 56.83 83.67 63.06 80.47 77.82 72.37
Ours 2.5 64 55.29 83.08 62.60 81.23 77.66 71.97
Ours* 2.5 128 56.48 83.04 63.57 81.56 78.77 72.68
Ours* 2.5 64 56.40 82.95 63.07 80.63 78.53 72.32
Table 10: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e Boolq Hella. Lamb. Mmlu Open. Piqa Truth. Wino. Avg.
16 16Bits 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
DL 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 47.35 79.63 82.60 56.10 69.53 62.00 33.60 77.69 32.80 72.06 61.34
4.5 DL 47.18 77.61 83.70 57.57 63.67 62.83 33.60 77.64 31.21 71.35 60.64
Ours 49.06 80.43 83.18 57.35 71.41 64.86 33.40 78.40 34.76 70.40 62.33
5 DL 50.43 79.17 84.10 58.22 68.62 64.77 34.60 78.62 34.52 72.61 62.57
Ours 49.83 80.93 84.19 57.92 72.19 65.43 35.80 79.22 35.37 71.03 63.19
6 DL 51.45 81.19 84.59 58.94 69.73 65.98 36.40 79.33 34.76 74.03 63.64
Ours 51.54 81.40 84.89 58.57 71.96 66.85 35.40 79.38 36.84 74.35 64.12
Table 11: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e Boolq Hella. Lamb. Mmlu Open. Piqa Truth. Wino. Avg.
16 16Bits 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.47 85.10 88.90 63.18 73.53 78.72 36.00 82.32 40.88 79.01 68.71
DL 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.90 85.82 88.53 63.63 75.45 79.97 37.00 81.88 40.76 80.27 69.32
4.5 DL 59.73 86.24 88.69 64.68 75.28 80.70 36.40 82.10 41.62 82.40 69.78
Ours 59.64 86.24 88.69 64.51 76.89 81.20 37.20 82.86 40.39 82.00 69.96
5 DL 60.15 86.57 88.84 64.94 75.65 81.08 38.20 82.70 41.13 81.77 70.10
Ours 61.18 86.95 89.17 64.89 77.18 81.23 37.40 83.19 41.00 80.90 70.31
6 DL 61.09 86.53 88.93 65.45 76.89 81.99 38.00 83.08 41.13 82.48 70.56
Ours 62.12 86.15 88.72 65.65 76.69 82.00 37.60 83.30 40.88 81.85 70.50
Table 12: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e Boolq Hella. Lamb. Mmlu Open. Piqa Truth. Wino. Avg.
16 16Bits 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
DL 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 53.67 80.18 85.08 57.99 66.27 68.30 33.60 76.66 43.08 68.90 63.37
4.5 DL 50.09 77.74 85.96 58.98 67.86 68.39 34.60 75.84 43.33 68.51 63.13
Ours 53.50 81.65 85.90 59.32 69.05 69.59 33.60 78.62 44.92 68.98 64.51
5 DL 51.11 80.26 86.85 59.54 68.27 69.42 35.20 76.77 44.31 68.67 64.04
Ours 54.61 82.62 85.90 59.78 69.67 70.43 33.60 78.07 45.90 69.77 65.04
6 DL 53.50 82.45 86.51 60.89 68.93 70.64 36.80 77.91 46.39 70.48 65.45
Ours 54.18 82.24 86.30 60.68 69.73 70.52 34.80 78.73 45.78 70.01 65.30
Table 13: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e Boolq Hella. Lamb. Mmlu Open. Piqa Truth. Wino. Avg.
16 16Bits 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
DL 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.84 80.98 85.84 53.09 61.91 68.88 32.00 75.30 36.35 70.72 61.89
4.5 DL 50.51 79.38 86.02 52.96 60.76 68.91 28.60 73.78 33.17 64.01 59.81
Ours 53.75 81.61 86.21 54.26 64.16 70.25 31.40 75.57 37.09 68.51 62.28
5 DL 51.11 80.30 86.09 54.23 62.04 70.26 31.40 75.19 35.62 68.27 61.45
Ours 54.10 81.82 86.42 54.62 64.20 71.53 30.80 76.01 37.58 68.27 62.54
6 DL 53.84 81.57 86.94 55.36 62.10 71.46 31.20 75.95 35.50 66.06 62.00
Ours 52.65 81.65 86.97 55.57 64.06 71.54 30.20 76.17 35.74 67.32 62.19
Table 14: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e Boolq Hella. Lamb. Mmlu Open. Piqa Truth. Wino. Avg.
16 16Bits 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 55.38 81.82 85.54 61.54 65.67 76.39 34.40 78.94 38.80 72.22 65.07
DL 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 59.47 83.16 87.19 61.70 68.78 78.39 35.40 79.71 41.74 73.01 66.86
4.5 DL 56.23 83.33 86.88 62.40 67.49 78.93 34.60 79.11 40.64 70.88 66.05
Ours 57.59 84.09 88.23 62.51 69.42 79.51 33.00 79.49 41.00 74.03 66.89
5 DL 56.83 83.38 84.50 63.18 66.25 79.58 34.00 79.92 40.39 71.98 66.00
Ours 57.85 84.47 87.92 63.33 67.22 80.10 35.60 80.25 40.51 74.43 67.17
6 DL 57.51 83.04 86.12 63.28 67.69 80.02 36.40 80.36 40.15 71.51 66.61
Ours 58.19 83.88 86.85 63.44 68.12 80.40 36.40 80.47 40.88 73.48 67.21