### v2

# SignRoundV2: Toward Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

Abstract
Extremely low-bit quantization is critical for efficiently deploying Large Language Models (LLMs), yet it often leads to severe performance degradation at 2 bits and even at 4 bits (e.g., MXFP4). We present SignRoundV2, a post-training quantization framework designed to maintain high performance even under aggressive compression. SignRoundV2 introduces (1) a simple yet efficient adaptive mixed-precision strategy that leverages gradient information and quantization-induced reconstruction errors to guide layer-wise bit allocation, and (2) a set of lightweight stabilization techniques, including loss filtering and a pre-tuning scale search, to improve tuning effectiveness in extremely low-bit regimes. Our approach takes a significant step toward closing the performance gap between quantized and full-precision models. Experimental results across diverse LLMs demonstrate that SignRoundV2 achieves near-lossless performance in mixed MXFP settings, narrowing the gap to $\sim$1% at an average of 4.5 bits, while substantially improving accuracy in challenging 2-bit weight-only quantization. The source code is available at https://github.com/intel/auto-round.
11footnotetext: These authors contributed equally.22footnotetext: Correspondence: wenhua.cheng@intel.com33footnotetext: Working done during the internship at Intel.
1 Introduction
The advent of Large Language Models (LLMs) such as GPT [8], LLaMA [60, 62], Qwen [4, 70] and DeepSeek [41] has marked a major shift in modern artificial intelligence. As parameter counts scale from billions to hundreds of billions, these models have demonstrated unprecedented capabilities in reasoning, coding, multimodal understanding, and autonomous agent behaviors. However, such scaling comes at the cost of dramatically increased memory consumption, bandwidth pressure, and inference latency, posing significant barriers for real-world deployment, especially in resource-constrained environments such as consumer GPUs, CPUs, and edge devices.
To alleviate these limitations, post-training quantization (PTQ) [69, 22, 39, 12] has emerged as one of the most practical approaches because it avoids costly re-training and can be applied to a wide range of pretrained LLMs. By compressing weights and activations to low-bit representations, PTQ yields substantial reductions in memory footprint and hardware cost. In particular, the push toward extremely low-bit quantization, for example, sub-4-bit weight-only quantization and 4-bit activation quantization, has become a crucial enabler for democratizing LLM deployment.
(a) W2A16 results on Llama 2/3 70B
(b) Recovery ratio at 2.5 bits
Figure 1: Left: results on pure 2 bits (W2A16) for Llama 2/3 70B. Right: recovery ratio for adaptive bit assignments at 2.5 bits across four models.
Recent research has moved beyond conventional W4A16 and W8A8 schemes (where W and A denote weight and activation bit-widths) toward more expressive low-bit data types. For example, MXFP4 [1] and NVFP4 [54] are floating-point variants provided by modern accelerators, demonstrating that well-designed data types can maintain reasonable accuracy even under aggressive activation quantization. Despite these architectural advances, achieving high-accuracy recovery remains non-trivial, primarily due to the prevalent outliers in large-scale model. In parallel, weight-only quantization continues to push the limits of bit efficiency. In particular, BitNet [64, 68] explores quantization down to approximately 1.58 bits, showing that extremely low-bit representations can still be learned with specialized training or distillation. However, these methods require training from scratch or costly post-training distillation, making them less suitable for scenarios where fast, scalable, and hardware-efficient PTQ is desired.
To minimize the accuracy loss caused by quantization, adaptive mixed-precision quantization has emerged as a natural solution to the limitations of uniform quantization. By strategically allocating higher precision to sensitive layers while aggressively compressing more robust ones, it effectively balances model performance with computational efficiency [77, 17, 65, 73]. Therefore, the central challenge is the design of an accurate and efficient layer sensitivity metric to guide bit allocation. Early methods estimate sensitivity using Hessian-based metrics [17, 73, 53], or rely on layer-wise or model-wise output distortion measured on a small calibration dataset [30]. However, these heuristics often exhibit imperfect correlation when applied layer-wise, and model-wise approaches typically require slow calibration, especially for large language models (LLMs). In contrast, learning-based or reinforcement-learning-based approaches automate precision assignment [47, 65, 35], but at the cost of substantial compute overhead.
To address these challenges, we introduce SignRoundV2, an enhanced post-training quantization framework building upon SignRound [12]. Our method first utilizes a simple yet efficient adaptive mixed-precision strategy that integrates gradient information with quantization-induced deviations to capture both local parameter distortion and global task impact. By leveraging these reliable sensitivity signals, SignRoundV2 employs dynamic programming [5] to efficiently determine the optimal layer-wise bit configuration for a given budget, substantially narrowing the performance gap between quantized and full-precision models even under aggressive compression. Second, we emphasize that stabilizing the optimization process is important in extremely low-bit regimes. Building on findings that quantization parameter initialization is critical [39, 49], we introduce a lightweight pre-tuning search, inspired by the importance matrix in llama.cpp [25], to rapidly identify high-quality initial scales before the main tuning stage. To further ensure optimization stability, we incorporate a loss filtering mechanism that excludes loss outliers during tuning, preventing outlier gradients from degrading the learned quantization parameters. Figure 1 compares our method to state-of-the-art baselines, showing that it achieves similar performance at significantly lower cost and avoids the issues associated with Quantization-Aware Training (QAT), which are discussed in Section 2. In summary, our contributions are three-fold.
  • We introduce an efficient adaptive mixed-precision strategy that leverages gradient information and quantization-induced deviations to guide layer-wise bit allocation via dynamic programming, maintaining high performance even under aggressive compression.
  • We employ simple yet effective stabilization techniques, such as loss filtering and a lightweight pre-tuning scale search, to improve stability and accuracy in extremely low-bit regimes.
  • We demonstrate that SignRoundV2 achieves competitive accuracy compared with recent high-cost QAT and vector-quantization methods in weight-only quantization, while requiring substantially lower quantization cost and maintaining strong performance in weight-activation scenarios.
2 Related Work
Quantization-Aware Training (QAT).
QAT integrates learnable quantizers directly into the training loop, minimizing the task loss and fine-tuning additional parameters beyond the quantizers introduced by low-precision operations, as described in EfficientQAT [10], LLM-QAT [44], DL-QAT [32], and BitDistiller [18]. Recent works [9, 63, 20] have revisited vector quantization for extremely low-bit settings, but these methods still rely on QAT to achieve high accuracy. Despite their effectiveness, QAT methods suffer from several practical limitations. First, because they optimize the task loss and may update non-quantizer parameters, they can be more sensitive to the fine-tuning data distribution and may introduce forgetting or domain-specific overfitting [48, 31]. Second, QAT often requires careful hyperparameter tuning, and the practical quantization cost can be substantially higher than that of PTQ. Third, QAT typically relies on more training data than PTQ to achieve stable generalization, further increasing the overall resource requirements.
Post-Training Quantization (PTQ).
Post-training quantization (PTQ) avoids task-loss optimization and additional fine-tuning of model parameters and therefore offers a simple and resource-efficient pipeline for compressing LLMs [52, 46, 21, 71, 69, 66, 16]. Because LLM decoding is typically memory-bound, weight-only PTQ has become the dominant approach [22, 39, 11, 72, 3, 33, 12] in LLM inference. However, existing weight-only methods still suffer from a significant accuracy drop at extremely low bit-widths (e.g., 2-bit). Meanwhile, joint weight–activation quantization has emerged as a complementary direction, with techniques such as rotation-based transformations [9, 2, 45, 38, 58] proposed to address activation outliers, albeit at the cost of additional inference overhead.
Adaptive Mixed-Precision Quantization.
Adaptive mixed-precision quantization assigns heterogeneous bit-widths across layers or channels to better align with their varying sensitivity to quantization error. Early frameworks utilized second-order information, such as Hessian-based methods (e.g., HAWQ [17] and BAQ [75]), or reinforcement learning schemes like HAQ [65]. However, these approaches are prohibitively expensive for multi-billion-parameter LLMs due to the computational intensity of second-order matrix computations or extensive policy evaluations. While MixLLM [76] attempts to combine first- and second-order metrics for channel-level precision, its reliance on naive thresholding and neglect of activation distortion in weight-activation joint quantization often lead to suboptimal performance. To address these scalability issues, recent advances focus on optimizing the search space or eliminating search altogether. AMQ [35] treats mixed-precision assignment as an AutoML problem, employing search space pruning and quantization proxies to accelerate the process; nonetheless, it remains relatively slow for large-scale models. In contrast, SFMP [53] completely bypasses the computationally expensive search process by utilizing a fine-grained, hardware-friendly assignment logic. Similarly, MicroMix [43] applies FP4/FP6/FP8 precision at the channel level via MXFP microscaling kernels. While these methods achieve high efficiency, they often require specific hardware or specialized kernel support. In practical deployment, frameworks like llama.cpp [25] adopt lightweight heuristic strategies to assign mixed precision across different model architectures. Furthermore, given the rising prominence of Mixture-of-Experts (MoE) architectures, recent works such as MxMoE [19], MoQAE [59], and MoPEQ [13] have extended adaptive mixed-bit settings specifically to MoE models.
Rounding and Optimized Quantizer Search.
Rounding critically affects the quality of quantized weights. AdaRound [51] formulates weight rounding as a per-weight binary optimization problem, where a second-order Taylor expansion of the task loss approximates quantization-induced perturbations, and a layer-wise local reconstruction loss is minimized via continuous relaxation. This formulation has inspired subsequent works such as BRECQ [36] and SignRound [12]. FlexRound [34] increases rounding flexibility through element-wise scaling, albeit at the cost of introducing numerous hyperparameters. Oscillation-Free training [42] highlights instability arising from learnable rounding parameters. Activation-dependent rounding in AQuant [37] reduces activation error but is not applicable to weight-only inference. SignRound [12] jointly optimizes rounding and weight clipping using sign-based gradient descent, achieving substantial improvements, particularly at extremely low bit-widths such as 2 bits.
3 Methodology
This section presents our methodology. We begin in Section 3.1 by introducing the sensitivity metric, DeltaLoss, which guides mixed-bit assignment. Section 3.2 then leverages DeltaLoss with dynamic programming to determine adaptive bit-widths for each layer. Section 3.3 describes the tuning strategy for quantization parameters, while Section 3.4 introduces our loss exclusion strategy. Finally, Section 3.5 summarizes the main hyperparameters and presents a practical trick for loss computation.
Before introducing our method, we briefly review the standard quantize-dequantize (QDQ) formulation and SignRoundV1 [12]. The QDQ operator for weights $\mathbf{W}$ is defined as follows:
$\mathbf{QDQ(W)}=s\cdot\text{clip}\Big(\Big\lfloor\frac{\mathbf{W}}{s}\Big\rceil,n,m\Big),\quad n,m\in\mathbb{N},$ (1)
where the rounding operation $\lfloor\cdot\rceil$ is typically performed using the Round-to-Nearest (RTN) method. The scale factor $s$ is defined as:
$s=\frac{\max(\mathbf{W})-\min(\mathbf{W})}{2^{\text{bit}}-1},$ (2)
and for simplicity, we ignore the zero point.
SignRoundV1 [12] introduces three trainable parameters $v$, $\alpha$ and $\beta$, to reduce the quantization error. The parameter $v$ enhances the rounding operation
$\mathbf{QDQ(W)}=s\cdot\text{clip}\Big(\Big\lfloor\frac{\mathbf{W}}{s}+v\Big\rceil,n,m\Big),\quad n,m\in\mathbb{N},$ (3)
while $\alpha$ and $\beta$ refine the scale and zero-point through
$s=\frac{\max(\mathbf{W})\cdot\alpha-\min(\mathbf{W})\cdot\beta}{2^{\text{bit}}-1}.$ (4)
(a) W2A16 layer-wise DeltaLoss sensitivity
(b) MXFP4 layer-wise DeltaLoss sensitivity
Figure 2: Layer-wise DeltaLoss sensitivity of Llama-3.1-8B-Instruct under W2A16 and MXFP4.
3.1 DeltaLoss Sensitivity Metric
Prior work [17, 73] largely ignores first-order gradient information and instead relies on second-order approximations, such as the Hessian or Fisher information matrix, to estimate sensitivity under the assumption that gradients are near zero at the local optimum. However, this assumption breaks down when quantization induces a large loss. Moreover, although second-order metrics capture local curvature, they share a fundamental limitation with purely gradient-based methods: both fail to reflect the actual impact of quantization on the loss and therefore provide only a relative trend rather than a reliable measure of the true loss.
To address this limitation, we estimate the loss increase induced by the QDQ operation using a first-order Taylor approximation. Let $\mathcal{L}_{\mathrm{f}}$ and $\mathcal{L}_{\mathrm{q}}$ denote the losses of the full-precision and quantized models, respectively. We define the quantization-induced loss increase as
$\Delta\mathcal{L}=\mathcal{L}_{\mathrm{q}}-\mathcal{L}_{\mathrm{f}}.$ (5)
For a layer with full-precision weight tensor $W_{\mathrm{f}}$ and activation tensor $A_{\mathrm{f}}$, and their quantized counterparts $W_{\mathrm{q}}$ and $A_{\mathrm{q}}$, we approximate the loss increase by a first-order Taylor expansion around the quantized point:
$\Delta\mathcal{L}\approx g_{aq}^{\top}(A_{\mathrm{q}}-A_{\mathrm{f}})+g_{wq}^{\top}(W_{\mathrm{q}}-W_{\mathrm{f}}),$ (6)
where the gradients $g_{aq}=\partial\mathcal{L}/\partial A_{\mathrm{q}}$ and $g_{wq}=\partial\mathcal{L}/\partial W_{\mathrm{q}}$ are computed using the quantized model on a small calibration set.
While Eq. 6 provides a signed first-order estimate, positive and negative element-wise contributions may cancel each other out when aggregated over tokens, channels, and calibration samples. We therefore instead use the magnitude of the gradient-weighted perturbation as a more stable sensitivity score:
$\mathrm{DeltaLoss}=\left\|g_{aq}\odot(A_{\mathrm{q}}-A_{\mathrm{f}})\right\|_{1}+\left\|g_{wq}\odot(W_{\mathrm{q}}-W_{\mathrm{f}})\right\|_{1},$ (7)
where $\odot$ denotes element-wise multiplication and $\|\cdot\|_{1}$ denotes summation over all elements. The final layer-wise score is averaged over calibration samples.
This metric combines two complementary factors: the magnitude of the quantization perturbation and the task-aware gradient sensitivity. Layers that exhibit both large quantization deviations and large gradients receive higher DeltaLoss scores and are therefore assigned higher precision under a fixed bit budget.
For joint weight-activation quantization, both weight and activation perturbations can be included as in Eq. 7. In practice, prior studies [69, 66] indicate that activation quantization often dominates the degradation. Therefore, for memory-efficient mixed weight-activation settings where weights and activations share the same bit-width, we use the activation-induced term as the default sensitivity score:
$\mathrm{DeltaLoss}^{\mathrm{WA}}=\left\|g_{aq}\odot(A_{\mathrm{q}}-A_{\mathrm{f}})\right\|_{1}.$ (8)
Figure 2 visualizes the layer-wise sensitivity measured by DeltaLoss. It is evident that sensitivity varies across different data types, making it prohibitively labor-intensive or suboptimal to rely on heuristic rules to manually assign quantization bit-widths for each scheme. Moreover, certain layers within a block, such as down_proj, exhibit consistently higher sensitivity, suggesting that traditional block-wise fallback strategies may be suboptimal.
3.2 Layer-wise Bit Allocation
We formulate layer-wise bit allocation as a discrete optimization problem. Let $n$ denote the total number of layers, and let $B=\{b_{1},\dots,b_{K}\}$ be the set of allowed bit-widths. For each layer $i\in\{1,\dots,n\}$ and bit-width $b\in B$, we introduce a binary variable $I_{i,b}\in\{0,1\}$, where $I_{i,b}=1$ indicates that layer $i$ is assigned bit-width $b$.
For simplicity, we assume identical bit-widths for both weights and activations, although this formulation can be extended to accommodate different bit-widths or data types across layers.
Let $\Delta L_{i}(b)$ denote the DeltaLoss when quantizing layer $i$ to $b$ bits. Given a target average bit-width $T$, the mixed-precision assignment can be written as:
$\min_{I_{i,b}}$ $\sum_{i=1}^{n}\sum_{b\in B}\Delta L_{i}(b)\,I_{i,b}$ (9)
s.t. $\sum_{b\in B}I_{i,b}=1,\quad\forall i=1,\dots,n,$
$\sum_{i=1}^{n}\sum_{b\in B}b\,I_{i,b}P_{i}\leq T\sum_{i}P_{i},$
$I_{i,b}\in\{0,1\},\quad\forall i=1,\dots,n,\;b\in B.$
where $P_{i}$ denotes the number of parameters in layer ii. This optimization problem can be solved via dynamic programming [5] or as an integer linear program [67, 6], both of which are well-established approaches. We omit the details here and refer the reader to the existing literature.
3.3 Tuning Quantization Parameters
In SignRoundV1 [12], all quantization parameters are initialized with trivial values, specifically, the weight clipping parameters $\alpha$ and $\beta$ are set to 1.0, and the rounding perturbation $v$ is initialized to 0. A better initialization could potentially improve performance, similar to the effect of careful initialization in training neural networks [28, 26]. Although initializing the rounding values is challenging, prior works [39, 3, 25] have proposed training-free methods to search for weight clipping by aligning layer outputs or optimizing certain objectives. While effective, these approaches require layer-wise forward passes, which can be slow, or rely on floating-point zero points for better accuracy.
Inspired primarily by the importance matrix in llama.cpp [25], we adopt a simple variant to initialize the scale. The search process is formulated as
$\min_{s\in S}\;\frac{1}{N}\sum_{i=1}^{N}\left((W_{\mathrm{f}}-W_{\mathrm{q}})\odot\overline{A}\right)_{i}^{2},$ (10)
where $S$ denotes a predefined set of candidate scales, and $\overline{A}$ represents the channel-wise average of input activations, calibrated from a set of samples.
For symmetric quantization, which is the default setting used in this work, the set $S=\{s_{0},\ldots,s_{c}\}$ is constructed from the following candidates:
$s_{i}=\frac{\max(|\mathbf{W}|)}{2^{b-1}+\epsilon_{i}},\quad\epsilon_{i}\in\{-t,-t+\delta,\ldots,t\}.$ (11)
$t$ and $\delta$ are fixed for each scheme. Specifically, we set $t=0.9$ and $\delta=0.01$ for W2A16. We refine the optimal scale $s_{init}$ with a learnable parameter $\alpha$, i.e., $s=s_{init}\cdot\alpha$, where $\alpha\in[0,2.0]$. Finally, we follow SignRoundV1 [12] to tune $v$ in Equation 3.
3.4 Tuning Stabilization via Loss Filtering
To enhance tuning stability, we exclude the top-$k$ largest losses in a batch when computing the mean squared error between the quantized block output and the full-precision block, where $k$ is set to 0.1% of the total number of elements. Let $\mathrm{TopK}(\mathcal{L},k)$ denote the set of the $k$ largest losses in $\mathcal{L}$. The effective loss used for optimization is defined as follows, where $\setminus$” denotes the exclusion operation:
$L_{\text{eff}}=\frac{1}{|\mathcal{L}\setminus\mathrm{TopK}(\mathcal{L},k)|}\sum_{L_{i}\in\mathcal{L}\setminus\mathrm{TopK}(\mathcal{L},k)}L_{i}.$ (12)
3.5 Hyperparameters
For DeltaLoss, we use only 16 calibration samples with a sequence length of 256. For tuning, we follow the setup in SignRoundV1 [12]. Each transformer block is optimized for 200 steps using signed gradient descent, with a learning rate of $1/\text{steps}$ and a batch size of 8. The sequence length is fixed at 2048. To reduce quantization cost, we lower the default number of calibration samples from 512 to 128. We additionally provide a higher-cost recipe for improved accuracy, denoted as Ours*, where the number of steps is increased from 200 to 500, the learning rate is set to $2.0/\text{steps}$ and the number of calibration samples is restored to 512. Automatic mixed precision (AMP) is applied throughout to improve tuning efficiency.
Table 1: Average accuracies across five tasks (Section 4.2) using W2A16 at 2 bits and mixed W2A16 and W4A16G128 at 2.5 bits. “G” denotes the group size, and “Avg. Bits” denotes the average per-weight bit-width. Ours* indicates our improved recipe (Section 3.5). “-” denotes missing data from the original work.
Avg. Bits Method Group Size Llama2-7B Llama2-13B Llama2-70B Llama3-8B Llama3-70B
16 16-bit - 64.66 67.44 72.41 68.64 75.28
2 AQLM 2x8 57.61 62.22 69.85 - -
AQLM 1x16 61.85 64.95 70.84 64.10 70.10
QuIP# - 60.61 64.44 70.91 - -
EQAT 128 59.50 63.88 68.93 59.37 67.57
EQAT 64 60.14 63.48 69.48 60.76 67.89
GPTQ 128 41.56 48.29 34.38 - -
AWQ 128 34.74 35.99 35.49 - -
OmniQ 128 46.98 53.56 54.87 52.66 60.06
SRV1 128 54.50 60.72 67.70 55.25 64.76
Ours 128 57.88 61.88 68.39 57.90 69.02
Ours* 128 58.67 62.34 68.82 59.97 70.16
Ours* 64 59.04 62.81 69.30 60.38 70.93
2.5 Ours 128 59.96 63.65 70.10 62.76 72.37
Ours* 128 60.28 63.73 70.04 62.60 72.68
Ours 64 60.08 64.09 70.25 63.67 71.97
Ours* 64 60.44 64.57 70.60 63.56 72.32
4 Experiments
4.1 Experimental Setup
Models and Benchmarks.
We evaluate SignRoundV2 on two prominent LLM families: LLaMA [61, 27] and Qwen [4, 70]. Our evaluation spans both low-bit and mixed-precision quantization settings across a comprehensive suite of standard benchmarks. All evaluations are conducted in a zero-shot setting using the lm-evaluation-harness framework [24]. To ensure fair comparisons, all methods within each experimental group are evaluated with a version of lm_eval that produces identical or closely aligned baseline results, unless explicitly stated otherwise. The benchmark suite includes ARC-Challenge [15], ARC-Easy [15], BoolQ [14], HellaSwag [74], LAMBADA [55], MMLU [29], OpenBookQA [50], PIQA [7], TruthfulQA [40], and WinoGrande [56].
Quantization Configurations.
Following SignRoundV1 [12], we utilize the Pile [23] as the calibration dataset. Unlike SignRoundV1, we adopt symmetric quantization to enhance hardware and kernel compatibility. All experiments are executed on NVIDIA A100 (80GB) GPUs. For clarity, the bit costs associated with scales and zero points are omitted unless otherwise specified. For MXFP4 and MXFP8 settings, we strictly adhere to the standard definitions specified in [1].
4.2 Comparison of Weight-Only Quantization Methods
Table 1 reports average accuracies across five representative tasks for W2A16 with and without mixed-precision support. In uniform W2A16 settings, SignRoundV2 consistently outperforms established PTQ baselines, including GPTQ [22], AWQ [39], OmniQuant [57], and SignRoundV1 [12], by a substantial margin. Compared to high-complexity methods such as AQLM [20], QuIP# [63], and EfficientQAT [10], our approach achieves competitive performance on large-scale models while maintaining a significantly lower optimization overhead. Furthermore, at slightly higher average bit-widths (e.g., 2.5 bits), our adaptive layer-wise strategy effectively preserves model quality, demonstrating robust recovery capabilities in extreme compression regimes. Detailed INT2/4 mixed-bit results are in Appendix E.
4.3 Comparison of Adaptive Bit-Width Quantization
Table 2 shows that SignRoundV2 achieves strong accuracy across evaluated bit-widths, with particularly large gains in the ultra-low-bit regime and competitive performance at higher bit-widths. Following SFMP [53], we report average accuracy across six tasks, prioritizing acc_norm. To ensure fair comparison, we account for scaling factor and zero-point overhead while restricting the bit-width search space to a maximum difference of 1. For reference, we provide the accuracy of uniform-bit SignRoundV2 at 3 and 4 bits (shown in gray), serving as non-mixed baselines. Notably, SFMP [53] and AMQ [35] use AWQ [39] as their base weight quantizer. Against these baselines, SignRoundV2 achieves significant gains in ultra-low-bit regimes; e.g., on Llama-3.1-8B at 2.5 bits, it improves the recovery rate over AMQ by 12.24 percentage points. SignRoundV2 achieves competitive performance with SFMP while maintaining hardware-friendly per-tensor granularity, avoiding kernel-level overhead inherent in SFMP’s per-group strategy. These results, supported by Table 6, highlight our favorable accuracy–cost trade-off. We further compare DeltaLoss against naive strategies such as head/tail fallback in Appendix A.
4.4 Performance under MXFP Quantization
Table 3 presents our results using the MX format [1]. Given that most existing PTQ methods do not support MX formats or lack comparable baselines, we primarily compare SignRoundV2 against SignRoundV1 [12] and RTN. SignRoundV2 exhibits performance gains in the 4-bit regime, surpassing RTN and outperforming SignRoundV1 in most scenarios. At average bit-widths of 4.5 and 5 bits, our method approaches full-precision performance on most evaluated models. Detailed results for MXFP4/8 are deferred to Appendix C.2, while Appendix D reports the NVFP4 results.
Table 2: Comparison of mixed-bit algorithms. Average accuracy (acc_norm preferred) and recovery rate across six tasks (see Appendix B for details). Gray rows: uniform 3-bit and 4-bit baselines quantized with SignRoundV2. Note: Ours employs hardware-friendly per-tensor mixed-bit granularity, avoiding per-group granularity overhead in SFMP.
Avg. bits Method Llama3.1-8B Qwen3-8B Qwen3-14B Qwen3-32B
16 Ours 75.83 74.07 77.41 78.05
SFMP 75.01 74.20 77.38 77.99
2.5 AMQ 58.65 (78.19%) 58.62 (79.00%) 66.56 (86.02%) 69.37 (88.95%)
SFMP 64.34 (85.78%) 66.16 (89.16%) 71.12 (91.91%) 73.33 (94.02%)
Ours 68.57 (90.43%) 66.82 (90.20%) 72.63 (93.83%) 74.96 (96.04%)
3 AMQ 68.78 (91.69%) 68.40 (92.18%) 71.32 (92.17%) 74.02 (94.91%)
SFMP 69.74 (92.89%) 71.36 (96.17%) 75.30 (97.31%) 76.24 (97.76%)
Ours 71.64 (94.48%) 72.21 (97.49%) 74.66 (96.45%) 76.61 (98.16%)
W3G128 72.84 (96.06%) 73.24 (98.88%) 76.39 (98.68%) 76.66 (98.22%)
3.5 AMQ 72.56 (96.73%) 71.65 (96.56%) 75.76 (97.91%) 76.01 (97.46%)
SFMP 72.97 (97.28%) 72.74 (98.03%) 76.89 (99.37%) 77.24 (99.04%)
Ours 73.94 (97.51%) 72.99 (98.54%) 76.32 (98.59%) 77.36 (99.12%)
4 AMQ 73.46 (97.93%) 72.64 (97.90%) 76.62 (99.02%) 77.15 (98.92%)
SFMP 74.33 (99.09%) 73.29 (98.77%) 77.13 (99.68%) 77.89 (99.87%)
Ours 75.07 (99.00%) 73.76 (99.58%) 77.05 (99.53%) 77.48 (99.27%)
W4G128 75.06 (98.99%) 74.11 (100.05%) 77.33 (99.90%) 77.75 (99.62%)
Table 3: Mixed-precision MXFP 4/6-bit quantization results. Average accuracies and recovery rates (%) across 10 tasks. “Avg. Bits”: average bit-width per weight; “DL”: DeltaLoss without tuning; “I”: Instruct models. See Appendix C.1 for details.
Avg. Bits Method Llama3.1-8B-I Llama3.1-70B-I Qwen2.5-7B-I Qwen3-8B Qwen3-32B
16 16-bit 64.16 70.00 65.67 63.24 67.00
4 RTN 58.31 (90.88%) 68.76 (98.23%) 60.62 (92.32%) 58.54 (92.57%) 64.87 (96.82%)
SRV1 60.72 (94.64%) 69.01 (98.60%) 64.06 (97.55%) 60.25 (95.28%) 66.92 (99.88%)
Ours 61.47 (95.81%) 69.31 (99.01%) 63.72 (97.03%) 61.45 (97.17%) 65.90 (98.36%)
4.5 DL 62.00 (96.63%) 70.17 (100.24%) 64.16 (97.70%) 61.35 (97.01%) 66.26 (98.90%)
Ours 63.34 (98.72%) 70.31 (100.44%) 64.77 (98.63%) 62.74 (99.21%) 67.10 (100.15%)
5 DL 63.53 (99.02%) 70.69 (100.99%) 65.22 (99.31%) 62.06 (98.13%) 66.41 (99.12%)
Ours 64.00 (99.75%) 70.66 (100.94%) 65.35 (99.51%) 62.81 (99.32%) 67.19 (100.28%)
4.5 Ablation Studies
We conduct ablation studies to isolate the individual contributions of pre-tuning initialization and loss filtering. As summarized in Table 4, both strategies consistently enhance performance under the W2A16G64 setting. Notably, incorporating loss filtering yields a significant boost in average accuracy (+0.82% on Qwen3), while its synergy with proper initialization (Ours) achieves a substantial 3.41% absolute improvement on Llama3.1. These results underscore that stabilizing the initial optimization phase is paramount for preserving the reasoning capabilities of ultra-low-bit quantized models.
Table 4: Ablation studies of loss filtering and pre-tuning initialization at W2A16G64.
Model Method ARC-c ARC-e BoolQ Hella LAMB MMLU OBQA PIQA Truth. Wino AVG.
Qwen3-8B V1 38.65 73.11 80.83 44.60 48.28 56.19 28.40 71.82 30.84 63.14 53.59
V1+Filtering 45.14 75.97 78.62 44.78 51.50 55.23 25.80 72.36 31.21 64.48 54.51
V1+Init 40.78 73.82 76.94 44.75 51.29 56.77 28.60 71.33 32.07 64.25 54.06
Ours 43.52 75.13 74.89 45.10 52.80 57.56 26.00 71.22 31.70 64.64 54.26
Llama3.1-8B-I V1 36.95 69.82 73.06 44.89 50.15 43.80 25.20 73.18 29.74 65.19 51.20
V1+Filtering 36.86 69.99 70.89 46.72 55.48 46.05 27.80 72.31 31.58 65.75 52.34
V1+Init 36.52 71.76 80.21 46.87 59.11 48.28 26.40 72.91 29.74 66.69 53.85
Ours 38.65 71.21 77.46 47.27 60.45 49.67 26.20 73.61 29.87 67.48 54.19
4.6 Quantization Cost
Table 6 compares end-to-end quantization time. QAT-based and vector quantization methods (e.g., EfficientQAT, QuIP#, AQLM) typically require tens to hundreds of GPU hours and careful hyperparameter tuning, often leading to higher practical costs than reported. In contrast, SignRoundV1 completes in 2.2 hours, while our method requires 2.5 hours, and 6 hours for the enhanced version (Ours*). Overall, our approach achieves competitive accuracy with substantially lower time cost.
Table 6 reports the additional VRAM and runtime overhead of DeltaLoss for weight-only quantization (e.g., W2A16), which remains comparable to MXFP4.
Table 5: Runtime Cost Comparison for Llama2-70B
Method Fits A100-80GB GPU hours
SignRoundV1 $\checkmark$ 2.2
EfficientQAT $\checkmark$ 41
QuIP# $\times$ 270
AQLM $\checkmark$ 336
Ours $\checkmark$ 2.5
Ours* $\checkmark$ 6
Table 6: Cost Comparison for Adaptive Bit-Width Assignment (Excluding Quantization Overhead).
Model Method VRAM (GB) GPU hours
Qwen3-8B AMQ 20 7
SFMP 48 0.05
Ours 15 $0.017{\times}\text{len(opt)}$
Qwen3-32B AMQ 70 $\approx 48$
SFMP 150 $\approx 0.2$
Ours 30 $0.05{\times}\text{len(opt)}$
Llama2-70B AMQ 160 176
SFMP 350 0.6
Ours 40 $0.117{\times}\text{len(opt)}$
5 Conclusion
We present SignRoundV2, a robust post-training quantization framework engineered for extremely low-bit LLM deployment. By integrating a gradient-informed sensitivity metric with a lightweight pre-tuning mechanism, SignRoundV2 ensures stable and accurate quantization under aggressive compression regimes. Extensive experiments demonstrate that our approach maintains high-fidelity performance at 4–5 bits and yields robust results even at 2 bits. These results confirm that effective PTQ can significantly reduce the computational and memory footprints of LLMs, offering a scalable solution for efficient large-scale model inference.
6 Limitations
Despite its strong performance, SignRoundV2 has several limitations. First, in uniform 2-bit settings, it still exhibits a noticeable accuracy gap compared with full-precision baselines, indicating that extremely low-bit PTQ remains challenging without mixed precision or additional training. Second, the bit-width configuration is determined before block-wise tuning and remains static during optimization; jointly updating bit allocation and rounding parameters may further improve accuracy but would increase search complexity. Third, DeltaLoss relies on gradient computation, which limits its immediate applicability in inference-only frameworks such as ONNX Runtime.
Ethics Statement
This research aims to advance efficient LLM quantization. SignRoundV2 leverages open-source models and publicly available datasets, adhering to their respective licenses and usage terms. As our method requires only minimal tuning of pre-existing models and is not tied to specific high-risk applications, it carries no significant new ethical risks. We acknowledge the contributions of the creators and maintainers of these resources and provide proper citations to the original sources.
Appendix A Comparison with Naive Mixed-bit Methods
The results in Tables 7 and 8 demonstrate that our method consistently outperforms simple head-layer and tail-layer heuristics across a range of precision budgets and model architectures. For MXFP quantization (4.5, 5, and 6 bits), the DL strategy achieves the highest average accuracy for all three models without requiring any tuning, showing clear improvements over allocating 4-bit precision only to the head (near the LM head) or tail layers (near the embedding). At lower precision (average 3 bits using W2G128/W4G128), the performance gap becomes even more obvious: both head- and tail-focused heuristics degrade substantially, while our method remains robust and delivers the best accuracy across all evaluated models.
Table 7: Comparison with head-layer (Head) and tail-layer (Tail) heuristics at 4.5, 5, and 6 bits using DeltaLoss without tuning. The average accuracies are computed over 10 tasks; see Section 4.4 for details. The “I” in the model name is short for “Instruct”.
Avg. bits Method Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
4.5 Head 4-bit 58.93 61.38 59.26
Middle 8-bit 58.96 60.79 59.36
Middle 4-bit 58.94 62.21 59.50
Tail 4-bit 59.18 60.63 58.92
DL 60.64 63.13 59.81
5 Head 4-bit 60.12 61.35 59.65
Middle 8-bit 59.40 61.03 60.26
Middle 4-bit 59.61 62.79 60.20
Tail 4-bit 60.82 61.75 59.43
DL 62.57 64.04 61.45
6 Head 4-bit 62.65 63.39 61.53
Middle 8-bit 60.92 62.87 61.42
Middle 4-bit 61.80 62.91 60.65
Tail 4-bit 59.88 62.22 60.54
DL 63.64 65.45 62.00
Table 8: Average accuracies across 10 tasks at an average of 3 bits (W2G128 / W4G128). The “I” in the model name is short for “Instruct”.
Avg. bits = 3 Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
Head 4-bit 31.98 32.70 31.96
Tail 4-bit 60.58 37.98 45.36
Ours 61.48 40.58 48.62
Appendix B Comparison with Prior Mixed-bit Methods
This appendix provides the detailed six-task accuracies for the mixed-bit weight-only quantization results reported in Section 4.3. We report HellaSwag, WinoGrande, ARC-Easy, ARC-Challenge, PIQA, BoolQ, and the average accuracy (AVG) for each model and method. Gray rows denote uniform-bit baselines, namely W3G128 and W4G128. Specifically, Table 9 reports results for Llama-3.1-8B, Table 10 for Qwen3-8B, Table 11 for Qwen3-14B, and Table 12 for Qwen3-32B.
Table 9: Detailed accuracies on six tasks for Llama3.1-8B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Gray rows denote uniform-bit baselines.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 54.78 82.70 83.12 79.26 80.90 74.19 75.83
SFMP-16-bit 53.41 81.19 82.15 78.99 81.39 72.93 75.01
2.5 AMQ 34.89 59.63 65.57 57.18 71.00 63.61 58.65
SFMP 41.13 66.58 73.49 64.35 74.05 66.46 64.34
Ours 45.90 73.40 78.38 69.02 75.57 69.14 68.57
3 AMQ 45.48 72.69 76.48 70.38 77.64 70.01 68.78
SFMP 47.40 75.80 77.43 71.89 78.13 67.80 69.74
Ours 49.06 76.60 79.88 74.75 78.18 71.35 71.64
W3G128 51.88 78.79 81.44 75.87 78.18 70.88 72.84
3.5 AMQ 49.57 77.10 80.00 76.15 79.54 73.01 72.56
SFMP 49.49 76.72 81.16 76.95 79.87 73.64 72.97
Ours 52.82 80.18 81.35 76.94 79.16 73.16 73.94
4 AMQ 50.68 78.20 81.04 77.83 79.92 73.09 73.46
SFMP 52.22 79.50 81.53 77.76 81.01 73.95 74.33
Ours 54.18 81.82 81.93 78.38 79.60 74.51 75.07
W4G128 54.01 81.65 82.42 78.84 80.03 73.40 75.06
Table 10: Detailed accuracies on six tasks for Qwen3-8B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Grayed rows denote uniform 3-bit and 4-bit baselines quantized with SignRoundV2.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 56.14 80.77 86.61 74.98 77.97 67.96 74.07
SFMP-16-bit 56.65 80.85 86.64 74.93 77.47 68.66 74.20
2.5 AMQ 35.75 56.19 75.90 55.76 69.70 58.41 58.62
SFMP 44.37 72.05 83.09 61.46 72.69 63.30 66.16
Ours 44.97 72.81 81.77 62.30 73.29 65.75 66.82
3 AMQ 47.61 73.78 84.40 66.71 73.94 63.93 68.40
SFMP 54.18 79.08 84.56 70.10 75.24 65.00 71.36
Ours 53.58 78.70 86.12 68.74 76.61 69.53 72.21
W3G128 55.55 80.30 85.72 71.50 77.69 68.67 73.24
3.5 AMQ 51.02 77.06 86.40 71.42 76.93 67.08 71.65
SFMP 55.12 78.28 85.69 72.70 76.12 68.51 72.74
Ours 54.10 80.05 86.39 71.93 77.26 68.19 72.99
4 AMQ 53.92 78.49 85.29 73.64 77.25 67.27 72.64
SFMP 55.15 79.04 85.88 74.20 77.09 68.35 73.29
Ours 56.31 79.67 86.61 73.77 77.53 68.67 73.76
W4G128 57.42 80.26 86.88 73.79 77.26 69.06 74.11
Table 11: Detailed accuracies on six tasks for Qwen3-14B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Grayed rows denote uniform 3-bit and 4-bit baselines quantized with SignRoundV2.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 60.49 82.74 89.39 78.80 80.03 73.01 77.41
SFMP-16-bit 60.41 82.79 89.33 78.92 79.98 72.84 77.38
2.5 AMQ 44.18 70.47 84.39 64.31 72.09 63.90 66.56
SFMP 50.68 75.21 86.73 69.14 76.44 68.51 71.12
Ours 51.62 78.58 87.98 70.21 77.09 70.32 72.63
3 AMQ 50.27 75.89 85.33 71.16 75.94 69.34 71.32
SFMP 57.58 80.89 87.80 75.35 78.45 71.74 75.30
Ours 57.08 80.26 87.37 74.15 77.58 71.51 74.66
W3G128 59.56 82.15 88.01 76.12 79.27 73.24 76.39
3.5 AMQ 58.31 81.56 87.56 76.04 79.12 71.98 75.76
SFMP 59.98 82.49 88.96 77.35 79.60 72.93 76.89
Ours 59.04 81.94 88.99 77.05 79.22 71.67 76.32
4 AMQ 59.42 82.05 88.76 77.68 79.65 72.13 76.62
SFMP 60.41 83.08 89.02 78.23 79.60 72.45 77.13
Ours 59.81 82.58 88.90 78.19 79.71 73.09 77.05
W4G128 61.01 82.91 89.08 78.17 80.09 72.69 77.33
Table 12: Detailed accuracies on six tasks for Qwen3-32B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Gray rows denote uniform-bit baselines.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 60.92 83.29 86.42 82.60 81.99 73.09 78.05
SFMP-16-bit 60.92 83.25 86.42 82.56 81.88 72.93 77.99
2.5 AMQ 50.63 73.90 80.08 71.68 75.74 64.19 69.37
SFMP 55.89 78.41 82.26 76.93 79.16 67.32 73.33
Ours 55.97 79.21 88.41 74.50 78.24 73.40 74.96
3 AMQ 58.83 79.62 83.26 77.10 77.14 68.15 74.02
SFMP 59.64 81.35 86.70 80.00 79.27 70.48 76.24
Ours 58.79 80.81 89.02 78.29 79.49 73.24 76.61
W3G128 60.32 82.74 85.20 78.66 79.92 73.09 76.66
3.5 AMQ 59.71 81.15 84.78 80.02 79.14 71.26 76.01
SFMP 60.41 82.02 85.88 81.18 81.42 72.53 77.24
Ours 62.37 82.91 85.14 80.57 80.58 72.61 77.36
4 AMQ 60.87 82.31 85.42 81.58 80.95 71.76 77.15
SFMP 61.09 83.46 86.20 82.01 81.73 72.83 77.89
Ours 61.01 82.11 85.05 80.98 81.39 74.35 77.48
W4G128 61.43 82.37 85.75 82.24 81.28 73.40 77.75
Appendix C Mixed MXFP Results
C.1 Mixed MXFP4/6 Results
This subsection reports the detailed results for mixed MXFP4/6 quantization. We list the task-wise accuracies across 10 tasks for each model provided in Tables 13, 14, 15, 16, and 17.
Table 13: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
SRV1 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 50.68 80.05 83.12 56.28 69.05 61.69 32.40 78.13 32.19 71.11 61.47
4.5 DL 48.38 79.29 83.88 58.31 68.48 64.51 31.60 78.78 35.01 71.74 62.00
Ours 50.43 80.13 84.86 57.83 71.36 65.75 34.60 79.43 36.60 72.45 63.34
5 DL 50.94 81.06 85.11 58.90 68.64 66.47 35.00 79.54 35.25 74.35 63.53
Ours 51.11 81.52 84.68 58.77 71.76 66.81 34.60 79.43 36.35 74.98 64.00
Table 14: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.56 85.10 88.81 63.16 73.36 79.03 36.20 82.10 40.88 79.40 68.76
SRV1 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.39 85.02 89.27 63.12 74.75 79.92 37.60 82.21 40.64 81.14 69.31
4.5 DL 61.01 86.32 88.96 65.09 75.53 81.01 37.40 83.03 41.86 81.53 70.17
Ours 59.98 85.94 89.24 65.10 76.98 81.21 36.80 83.03 41.00 83.82 70.31
5 DL 61.86 86.57 88.81 65.57 76.79 81.75 37.40 83.35 42.59 82.16 70.69
Ours 61.35 86.36 88.75 65.89 76.98 82.19 38.00 83.35 40.76 82.95 70.66
Table 15: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
SRV1 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 51.28 80.81 85.72 58.16 66.47 68.49 34.20 77.64 45.29 69.14 63.72
4.5 DL 50.60 80.13 86.51 59.27 69.22 69.73 33.80 76.66 45.53 70.17 64.16
Ours 53.92 81.73 86.21 59.56 69.03 70.43 34.60 77.20 45.41 69.61 64.77
5 DL 52.47 81.78 86.73 60.66 68.41 70.32 36.80 77.69 46.39 70.96 65.22
Ours 53.50 82.07 86.33 60.75 69.36 70.77 35.00 78.07 46.88 70.80 65.35
Table 16: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
SRV1 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.92 81.65 85.50 52.75 62.14 69.41 30.40 75.19 35.62 67.88 61.45
4.5 DL 52.05 80.30 85.96 53.96 62.57 70.99 31.00 74.32 34.27 68.03 61.35
Ours 54.95 82.66 86.67 54.82 64.37 71.52 30.80 76.12 37.33 68.19 62.74
5 DL 51.54 80.22 87.25 55.43 62.41 71.96 31.80 76.17 35.62 68.19 62.06
Ours 54.44 82.53 86.97 55.75 64.27 71.90 31.80 76.33 35.62 68.51 62.81
Table 17: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 56.66 81.23 84.95 61.73 65.96 76.59 33.80 77.97 39.17 70.64 64.87
SRV1 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 55.97 82.70 85.87 61.23 68.79 78.41 34.60 78.62 40.15 72.69 65.90
4.5 DL 57.08 83.25 86.39 63.07 66.08 79.48 33.40 80.96 42.11 70.80 66.26
Ours 56.57 84.22 88.04 63.44 68.33 79.91 36.00 80.14 40.88 73.48 67.10
5 DL 57.85 84.22 85.20 63.35 67.53 79.84 35.60 79.27 39.53 71.67 66.41
Ours 60.15 84.22 87.25 63.38 68.45 80.15 35.40 80.09 39.53 73.32 67.19
C.2 Mixed MXFP4/8 Results
This subsection reports the detailed results for mixed MXFP4/8 quantization. Table 18 reports the average accuracies and recovery rates across 10 tasks, for MXFP4/8 mixed-precision quantization. The results show that our method generally improves over RTN and achieves competitive or better accuracy than SignRoundV1 (SRV1) in the 4-bit setting. At higher average bit widths (4.5–6 bits), DeltaLoss-only (DL) and our method approach full-precision accuracy, with our method achieving the best recovery in most scenarios. Overall, these results demonstrate that our approach effectively preserves model performance under extreme low-bit mixed-precision quantization while minimizing accuracy loss. The corresponding task-wise accuracies are provided in Tables 19, 20, 21, 22, and 23.
Table 18: Average accuracies and recovery rates (%) across 10 tasks, with MXFP4 at 4 bits and mixed MXFP4/8 for the higher-bit settings. “Avg. Bits” indicates the average per-weight bit-width. “DL” denotes DeltaLoss without tuning. The “I” in the model name is short for “Instruct”.
Avg. Bits Method Llama3.1-8B-I Llama3.1-70B-I Qwen2.5-7B-I Qwen3-8B Qwen3-32B
16 16-bit 64.16 70.00 65.67 63.24 67.00
4 RTN 58.31 (90.88%) 68.71 (98.16%) 60.62 (92.32%) 58.54 (92.57%) 65.07 (97.12%)
SRV1 60.72 (94.64%) 69.01 (98.60%) 64.06 (97.55%) 60.25 (95.28%) 66.92 (99.88%)
Ours 61.34 (95.59%) 69.32 (99.04%) 63.37 (96.50%) 61.89 (97.87%) 66.86 (99.79%)
4.5 DL 60.64 (94.50%) 69.78 (99.70%) 63.13 (96.13%) 59.81 (94.58%) 66.05 (98.58%)
Ours 62.33 (97.14%) 69.96 (99.95%) 64.51 (98.24%) 62.28 (98.49%) 66.89 (99.83%)
5 DL 62.57 (97.51%) 70.10 (100.15%) 64.04 (97.52%) 61.45 (97.18%) 66.00 (98.51%)
Ours 63.19 (98.49%) 70.31 (100.45%) 65.04 (99.03%) 62.54 (98.89%) 67.17 (100.25%)
6 DL 63.64 (99.18%) 70.56 (100.80%) 65.45 (99.67%) 62.00 (98.04%) 66.61 (99.42%)
Ours 64.12 (99.93%) 70.50 (100.71%) 65.30 (99.43%) 62.19 (98.34%) 67.21 (100.32%)
Table 19: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
SRV1 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 47.35 79.63 82.60 56.10 69.53 62.00 33.60 77.69 32.80 72.06 61.34
4.5 DL 47.18 77.61 83.70 57.57 63.67 62.83 33.60 77.64 31.21 71.35 60.64
Ours 49.06 80.43 83.18 57.35 71.41 64.86 33.40 78.40 34.76 70.40 62.33
5 DL 50.43 79.17 84.10 58.22 68.62 64.77 34.60 78.62 34.52 72.61 62.57
Ours 49.83 80.93 84.19 57.92 72.19 65.43 35.80 79.22 35.37 71.03 63.19
6 DL 51.45 81.19 84.59 58.94 69.73 65.98 36.40 79.33 34.76 74.03 63.64
Ours 51.54 81.40 84.89 58.57 71.96 66.85 35.40 79.38 36.84 74.35 64.12
Table 20: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.47 85.10 88.90 63.18 73.53 78.72 36.00 82.32 40.88 79.01 68.71
SRV1 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.90 85.82 88.53 63.63 75.45 79.97 37.00 81.88 40.76 80.27 69.32
4.5 DL 59.73 86.24 88.69 64.68 75.28 80.70 36.40 82.10 41.62 82.40 69.78
Ours 59.64 86.24 88.69 64.51 76.89 81.20 37.20 82.86 40.39 82.00 69.96
5 DL 60.15 86.57 88.84 64.94 75.65 81.08 38.20 82.70 41.13 81.77 70.10
Ours 61.18 86.95 89.17 64.89 77.18 81.23 37.40 83.19 41.00 80.90 70.31
6 DL 61.09 86.53 88.93 65.45 76.89 81.99 38.00 83.08 41.13 82.48 70.56
Ours 62.12 86.15 88.72 65.65 76.69 82.00 37.60 83.30 40.88 81.85 70.50
Table 21: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
SRV1 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 53.67 80.18 85.08 57.99 66.27 68.30 33.60 76.66 43.08 68.90 63.37
4.5 DL 50.09 77.74 85.96 58.98 67.86 68.39 34.60 75.84 43.33 68.51 63.13
Ours 53.50 81.65 85.90 59.32 69.05 69.59 33.60 78.62 44.92 68.98 64.51
5 DL 51.11 80.26 86.85 59.54 68.27 69.42 35.20 76.77 44.31 68.67 64.04
Ours 54.61 82.62 85.90 59.78 69.67 70.43 33.60 78.07 45.90 69.77 65.04
6 DL 53.50 82.45 86.51 60.89 68.93 70.64 36.80 77.91 46.39 70.48 65.45
Ours 54.18 82.24 86.30 60.68 69.73 70.52 34.80 78.73 45.78 70.01 65.30
Table 22: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
SRV1 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.84 80.98 85.84 53.09 61.91 68.88 32.00 75.30 36.35 70.72 61.89
4.5 DL 50.51 79.38 86.02 52.96 60.76 68.91 28.60 73.78 33.17 64.01 59.81
Ours 53.75 81.61 86.21 54.26 64.16 70.25 31.40 75.57 37.09 68.51 62.28
5 DL 51.11 80.30 86.09 54.23 62.04 70.26 31.40 75.19 35.62 68.27 61.45
Ours 54.10 81.82 86.42 54.62 64.20 71.53 30.80 76.01 37.58 68.27 62.54
6 DL 53.84 81.57 86.94 55.36 62.10 71.46 31.20 75.95 35.50 66.06 62.00
Ours 52.65 81.65 86.97 55.57 64.06 71.54 30.20 76.17 35.74 67.32 62.19
Table 23: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
Avg. Bits Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. Avg.
16 16-bit 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 55.38 81.82 85.54 61.54 65.67 76.39 34.40 78.94 38.80 72.22 65.07
SRV1 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 59.47 83.16 87.19 61.70 68.78 78.39 35.40 79.71 41.74 73.01 66.86
4.5 DL 56.23 83.33 86.88 62.40 67.49 78.93 34.60 79.11 40.64 70.88 66.05
Ours 57.59 84.09 88.23 62.51 69.42 79.51 33.00 79.49 41.00 74.03 66.89
5 DL 56.83 83.38 84.50 63.18 66.25 79.58 34.00 79.92 40.39 71.98 66.00
Ours 57.85 84.47 87.92 63.33 67.22 80.10 35.60 80.25 40.51 74.43 67.17
6 DL 57.51 83.04 86.12 63.28 67.69 80.02 36.40 80.36 40.15 71.51 66.61
Ours 58.19 83.88 86.85 63.44 68.12 80.40 36.40 80.47 40.88 73.48 67.21
Appendix D NVFP4 Results
This section provides detailed NVFP4 accuracy results reported in Table 24. Our method achieves the best average accuracy on four out of five models and remains competitive on Qwen3-32B.
Table 24: Comparison of SignRoundV1 (SRV1), RTN, and our method under NVFP4 quantization. We report accuracy (%) across 10 tasks (listed in Section 4.1) and the average (AVG). The “I” in the model name is short for "Instruct".
Model Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. AVG
Llama3.1-8B-I 16-bit 53.75 82.15 85.35 59.79 72.06 68.28 35.60 80.30 38.07 73.32 64.87
RTN 50.60 80.01 83.43 58.30 71.34 64.13 35.00 78.89 34.27 71.90 62.79
SRV1 50.26 80.39 84.28 58.23 71.07 65.39 34.00 79.22 35.86 72.77 63.15
Ours 50.68 81.65 85.44 57.93 71.49 64.91 35.00 79.49 37.58 74.51 63.87
Llama3.1-70B-I 16-bit 61.09 86.20 89.11 66.28 76.46 82.52 38.80 83.73 41.98 83.43 70.96
RTN 60.58 85.94 89.45 65.03 75.47 80.54 37.20 83.24 42.11 81.14 70.07
SRV1 59.73 85.98 88.84 64.77 75.59 81.31 39.00 83.13 40.88 82.32 70.16
Ours 60.75 86.87 88.87 65.10 75.82 81.25 38.20 82.54 41.00 81.29 70.17
Qwen2.5-7B-I 16-bit 53.07 81.69 86.33 62.06 69.38 71.71 34.60 79.71 47.98 71.03 65.76
RTN 51.96 80.18 86.24 60.65 67.15 70.47 33.20 78.51 43.82 67.96 64.01
SRV1 52.82 81.14 86.09 60.90 69.07 70.61 34.80 78.67 47.61 70.40 65.21
Ours 54.18 81.69 86.79 60.23 69.65 70.67 34.00 78.51 47.12 70.96 65.38
Qwen3-8B 16-bit 55.20 83.54 86.61 57.16 64.10 72.90 31.60 76.66 36.47 67.96 63.22
RTN 53.41 82.37 85.87 55.29 62.45 70.82 30.80 76.44 33.66 67.56 61.87
SRV1 54.86 82.91 86.88 55.31 62.95 71.97 31.20 76.28 36.35 67.48 62.62
Ours 55.29 83.04 86.85 55.41 63.71 71.36 29.20 75.79 37.94 69.22 62.78
Qwen3-32B 16-bit 58.02 84.39 86.42 63.89 67.13 80.74 36.20 81.23 38.56 73.09 66.97
RTN 57.51 84.01 87.03 63.39 65.83 79.80 35.40 79.43 39.53 71.51 66.34
SRV1 59.56 84.22 88.01 62.99 68.29 79.50 37.60 79.54 41.00 72.22 67.29
Ours 57.51 84.68 86.51 63.09 68.89 79.80 36.00 79.60 39.05 72.38 66.75
Appendix E INT2/4 Mixed-bit Results
This appendix provides the detailed task-wise results for the INT2/4 mixed-bit weight-only quantization experiments discussed in Section 4.2. We report accuracies on ARC-Challenge, ARC-Easy, HellaSwag, PIQA, and WinoGrande for both uniform and mixed-precision settings across different model scales. These results complement the averaged comparisons in Table 1 and provide a more fine-grained view of the behavior of each method under extreme low-bit quantization. Specifically, Table 25 reports the detailed accuracies for Llama-3 models, and Table 26 reports the results for Llama-2 models.
Table 25: The detailed accuracies of 5 tasks for Llama3 models with INT2/4 mixed precision. “Avg. Bits” indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. PIQA Wino. Avg.
Llama3-8B 16-bit 16 - 50.17 80.09 60.14 79.54 73.24 68.64
AQLM 2 1x16 41.21 74.24 55.44 77.80 71.82 64.10
EfficientQAT 2 128 36.01 69.15 50.74 75.30 65.67 59.37
EfficientQAT 2 64 37.03 71.17 51.86 76.03 67.72 60.76
SignRoundV1 2 128 34.22 67.09 44.15 71.65 59.12 55.25
Ours 2 128 35.67 70.03 47.70 72.96 63.14 57.90
Ours 2 64 37.37 70.29 48.69 74.59 66.14 59.42
Ours* 2 128 38.99 72.18 49.19 73.94 65.57 59.97
Ours* 2 64 39.42 71.42 50.04 74.43 66.61 60.38
Ours 2.5 128 41.72 74.87 52.81 75.57 68.82 62.76
Ours 2.5 64 43.34 75.42 53.27 76.06 70.24 63.67
Ours* 2.5 128 41.38 74.87 53.31 75.46 67.96 62.60
Ours* 2.5 64 43.00 74.87 54.13 76.28 69.53 63.56
Llama3-70B 16-bit 16 - 60.24 86.83 66.40 82.21 80.74 75.28
AQLM 2 1x16 50.34 78.83 63.47 79.65 78.22 70.10
EfficientQAT 2 128 48.81 79.25 60.75 79.60 69.46 67.57
EfficientQAT 2 64 49.06 77.40 61.60 77.37 74.03 67.89
SignRoundV1 2 128 47.87 79.71 43.96 78.62 73.64 64.76
Ours 2 128 49.91 80.81 59.75 79.49 75.14 69.02
Ours 2 64 52.22 81.73 60.93 80.09 75.69 70.13
Ours* 2 128 52.05 82.41 60.41 79.38 76.56 70.16
Ours* 2 64 54.95 82.53 61.05 79.82 76.32 70.93
Ours 2.5 128 56.83 83.67 63.06 80.47 77.82 72.37
Ours 2.5 64 55.29 83.08 62.60 81.23 77.66 71.97
Ours* 2.5 128 56.48 83.04 63.57 81.56 78.77 72.68
Ours* 2.5 64 56.40 82.95 63.07 80.63 78.53 72.32
Table 26: The detailed accuracies of 5 tasks for Llama2 models with INT2/4 mixed precision. “Avg. Bits” indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. PIQA Wino. Avg.
Llama2-7B 16-bit 16 - 42.83 75.97 57.27 77.91 69.30 64.66
AQLM 2 2x8 32.85 66.92 49.96 73.07 65.27 57.61
AQLM 2 1x16 39.68 74.07 53.42 76.88 65.19 61.85
QuIP# 2 - 37.88 71.84 52.19 75.46 65.67 60.61
EfficientQAT 2 128 36.52 69.78 50.84 74.16 66.22 59.50
EfficientQAT 2 64 36.86 70.96 51.58 75.30 65.98 60.14
GPTQ 2 128 21.25 40.45 32.59 58.32 55.17 41.56
AWQ 2 128 21.08 24.62 25.69 52.34 49.96 34.74
OmniQ 2 128 23.46 50.13 40.28 65.13 55.88 46.98
SignRoundV1 2 128 32.25 65.99 40.28 72.96 61.01 54.50
Ours 2 128 33.53 69.32 48.17 74.43 63.93 57.88
Ours 2 64 32.93 67.97 48.54 74.21 65.59 57.85
Ours* 2 128 34.64 69.95 48.92 74.43 65.43 58.67
Ours* 2 64 35.32 69.99 49.30 74.16 66.41 59.04
Ours 2.5 128 35.41 71.46 50.62 75.30 67.01 59.96
Ours 2.5 64 35.92 71.25 51.44 75.35 66.46 60.08
Ours* 2.5 128 36.01 72.01 51.19 74.92 67.25 60.28
Ours* 2.5 64 36.09 71.72 51.85 74.81 67.72 60.44
Llama2-13B 16-bit 16 - 47.18 78.32 60.25 79.22 72.22 67.44
AQLM 2 2x8 40.10 73.06 54.62 77.09 66.22 62.22
AQLM 2 1x16 43.52 75.25 57.62 78.29 70.09 64.95
QuIP# 2 - 42.92 75.72 56.53 77.97 69.06 64.44
EfficientQAT 2 128 42.83 75.04 55.66 76.99 68.90 63.88
EfficientQAT 2 64 41.89 74.83 55.27 77.04 68.36 63.48
GPTQ 2 128 21.93 55.60 41.06 67.08 55.80 48.29
AWQ 2 128 23.12 26.22 25.80 52.99 51.85 35.99
OmniQ 2 128 30.29 63.22 46.23 70.13 57.93 53.56
SignRoundV1 2 128 38.57 71.17 53.35 76.17 64.33 60.72
Ours 2 128 39.08 73.78 52.96 75.68 67.88 61.88
Ours 2 64 40.53 74.33 53.87 76.44 68.43 62.72
Ours* 2 128 40.44 74.24 53.15 75.90 67.96 62.34
Ours* 2 64 40.78 74.58 53.73 76.77 68.19 62.81
Ours 2.5 128 41.04 75.42 55.39 76.77 69.61 63.65
Ours 2.5 64 41.13 75.55 55.81 77.42 70.56 64.09
Ours* 2.5 128 40.02 75.08 55.61 76.82 71.11 63.73
Ours* 2.5 64 42.83 76.43 55.97 77.37 70.24 64.57
Llama2-70B 16-bit 16 - 54.44 82.70 64.77 82.15 77.98 72.41
AQLM 2 2x8 51.45 79.76 61.94 80.47 75.61 69.85
AQLM 2 1x16 52.99 81.36 62.78 81.07 76.01 70.84
QuIP# 2 - 52.65 81.90 62.86 81.39 75.77 70.91
EfficientQAT 2 128 49.23 80.01 61.58 80.20 73.64 68.93
EfficientQAT 2 64 50.77 80.13 61.78 80.14 74.59 69.48
GPTQ 2 128 22.70 25.08 25.04 49.51 49.57 34.38
AWQ 2 128 22.35 25.76 25.46 52.50 51.38 35.49
OmniQ 2 128 33.28 67.21 35.45 74.10 64.33 54.87
SignRoundV1 2 128 46.59 78.37 59.65 79.00 74.90 67.70
Ours 2 128 49.06 79.59 59.99 78.35 74.98 68.39
Ours 2 64 49.06 78.91 60.21 79.00 75.61 68.56
Ours* 2 128 48.81 80.22 60.23 78.73 76.09 68.82
Ours* 2 64 50.60 79.92 60.52 79.22 76.24 69.30
Ours 2.5 128 51.71 80.56 61.77 79.82 76.64 70.10
Ours 2.5 64 50.77 80.77 61.91 80.30 77.51 70.25
Ours* 2.5 128 50.64 80.64 61.68 79.89 77.35 70.04
Ours* 2.5 64 51.02 81.31 62.29 80.58 77.82 70.60